### v1

# Fishers Harvest Parallel Unlearning in Inherited Model Networks

Abstract.
Unlearning in various learning frameworks remains challenging, with the continuous growth and updates of models exhibiting complex inheritance relationships. This paper presents a novel unlearning framework, which enables fully parallel unlearning among models exhibiting inheritance. A key enabler is the new Unified Model Inheritance Graph (UMIG), which captures the inheritance using a Directed Acyclic Graph (DAG). Central to our framework is the new Fisher Inheritance Unlearning (FIUn) algorithm, which utilizes the Fisher Information Matrix (FIM) from initial unlearning models to pinpoint impacted parameters in inherited models. By employing FIM, the FIUn method breaks the sequential dependencies among the models, facilitating simultaneous unlearning and reducing computational overhead. We further design to merge disparate FIMs into a single matrix, synchronizing updates across inherited models. Experiments confirm the effectiveness of our unlearning framework. For single-class tasks, it achieves complete unlearning with 0% accuracy for unlearned labels while maintaining 94.53% accuracy for retained labels on average. For multi-class tasks, the accuracy is 1.07% for unlearned labels and 84.77% for retained labels on average. Our framework accelerates unlearning by 99% compared to alternative methods.
1. Introduction
The explosive growth of data volume and the increasing complexity of applications necessitate models that can constantly update to adapt to new data, constraints, and standards (Jayaraman and Evans, 2019). Accordingly, the need for data privacy, the auditing of illegal information from models, and compliance with regulations and industry standards arises. This leads to the concept of “model forgetting” or “unlearning,” which involves removing specific information from trained models.
The importance of unlearning lies in its ability to remove data that may raise privacy concerns, ensuring that the model can no longer recognize or rely on such data, in compliance with regulations such as the General Data Protection Regulation (GDPR) (Voigt and Von dem Bussche, 2017). Additionally, unlearning plays a crucial role in data correction, removing the impact of erroneous or biased data from the model and enhancing its accuracy and fairness. This not only protects user privacy but also boosts the model’s quality and reliability, aligning with both ethical standards and legal requirements.
Machine unlearning becomes particularly challenging when models are consistently recorded with a directed acyclic graph (DAG) topology, where models inheriting from predecessors also require updates. Examples of such scenarios include a basic model being customized to meet various Machine Learning Operations (MLOps) lifecycle requirements (Matsui and Goya, 2023; Kreuzberger et al., 2023) or distilled to fit computational resource constraints (Polino et al., 2018).
The concept of model inheritance can be applied to prevalent learning frameworks, such as Federated Learning (FL) (Li et al., 2020a), Distributed Data-Parallel Learning (DDPL) (Li et al., 2020c), Incremental Learning (IL) (Van de Ven et al., 2022), and Transfer Learning (TL) (Weiss et al., 2016). In this case, unlearning even a single model would need to unlearn the entire subgraph rooted at the origin to ensure all descendant models are consistent and relevant.
Research Gap. Various machine unlearning techniques have been proposed. Unfortunately, the inheritance relationships among related models prevent most existing unlearning methods from being executed in parallel, including re-training (Kim and Woo, 2022; Liu et al., 2024b), gradient ascent (Zhang et al., 2024), or knowledge distillation methods (Sinha et al., 2023). On the other hand, Sharding, Isolation, Slicing, and Aggregated Training (SISA) (Bourtoule et al., 2021) is a commonly used efficient method that reduces the dependence between data by dividing it into multiple segments and training multiple models in parallel. However, SISA is also inapplicable when there are inheritance relationships between models (Liu et al., 2024a). This is due to the fact that SISA needs to be performed sequentially from the starting point of updating to the end of the subgraph, incurring significant overhead and resource consumption during unlearning. In this sense, there is a pressing need for an effective and efficient way to perform unlearning in the presence of model inheritance relationships.
Research Question. To address this research gap, we need to identify the commonalities across various prevalent frameworks in unlearning and determine how a new unlearning design can be universally applied. Specifically, we raise two key Research Questions (RQs):
RQ1. How can unlearning requests be swiftly allocated to models that need updates and then prepare these models for unlearning tasks within a large-scale, interdependent model network?
RQ2. How can the model unlearning tasks be efficiently and effectively executed, especially when the unlearning of one model impacts subsequent models?
In response to these two RQs, we investigate the topological structure and unlearning impact of inherited models in four prevalent frameworks: FL, DDPL, IL, and TL. Specifically, we study the key factors affecting the efficiency, accuracy, and consistency of model unlearning. These include the selection of model parameters, changes in data distribution, and utilization of computational resources.
Contributions. This paper presents a novel parallel unlearning framework that can conduct efficient unlearning among large numbers of dependent models with complex inheritance relations, e.g., in FL, DDPL, IL, and TL. We first propose a new Unified Model Inheritance Graph (UMIG), to unveil the inheritance characteristic among models in various learning frameworks. With the assistance of the UMIG and Fisher Information Matrix (FIM), we develop the Fisher Inheritance Unlearning (FIUn) method that can efficiently pinpoint critical model parameters to be unlearned.
This allows independent unlearning adjustments on large-scale inherited models. To further enable effective one-shot unlearning of knowledge inherited from multiple upstream, we design a merging FIM function that consolidates disparate FIMs and identifies the critical parameters that need updating across all sources.
The contributions of this paper are summarized as follows:
  • We propose the UMIG, which abstracts various prevalent learning frameworks, e.g., FL, DDPL, IL, and TL, in a unified fashion using DAGs to represent models’ interdependence. This facilitates the analysis of large-scale inheritance-oriented model networks.
  • We design the FIUn method, which enables efficient parallel unlearning within the UMIG. The method leverages the FIM to quantify the significance of model parameters to unlearning tasks, and selectively scales the parameters according to their importance.
  • We further develop the merging FIM function to consolidate the FIMs from multiple upstream unlearning models into a cohesive matrix, thereby aggregating the unlearning tasks inherited from those models. This design facilitates the one-shot removal of inherited knowledge and reduces computational overhead within the UMIG.
Comprehensive experiments demonstrate that our method achieves complete unlearning for single-class unlearning tasks across all considered learning frameworks and model types, retaining an average accuracy of 94.53%. For multi-class unlearning tasks, the average unlearning accuracy drops to 1.07%, while the accuracy for retained labels reaches up to 84.77%. Additionally, our algorithm shows a 99% speed improvement across all considered frameworks compared to alternative methods.
Paper Organization. The rest of the paper is organized as follows. Section 2 reviews the preliminaries. Section 3 provides the proposed UMIG, followed by introducing the novel FIUn method and merging FIM function in Section 4. Section 5 presents the experimental settings. Section 6 presents our comprehensive experimental results based on four learning frameworks. Section 7 reviews the related works. Section 8 concludes this work.
2. Preliminaries
This section introduces the concepts, notation, and formulas of this paper, including machine unlearning and FIM.
2.1. Machine Unlearning
Machine unlearning refers to removing the influence of specific data subsets from a trained model while maintaining its performance on the remaining data (Foster et al., 2024a; Umer and Polikar, 2021). In other words, the goal is to eliminate the impact of specific labels on the model’s predictions while ensuring that the prediction accuracy for the remaining labels and the overall performance of the model remain unaffected.
Let $D=\left\{\left(x_{i},y_{i}\right)\right\}^{N}_{i=1}$ represent a dataset where $x_{i}$ is the $i$-th training sample and $y_{i}\in\left\{1,\cdots,K\right\}$ is the corresponding class label. Let $C_{f}$ denote the set of labels to be unlearned, corresponding to the data set $D^{f}\subset D$ to be unlearned, and $C_{r}$ represent the set of retained labels, corresponding to the remaining data set $D^{r}=D\backslash D^{f}$.
Unlearning aims to eliminate the impact of $D^{f}$ with labels in $C_{f}$ from the model while preserving its performance on $D^{r}$ with labels in $C_{r}$. In this context, we define $\phi_{\theta}\left(\cdot\right):X\rightarrow Y$ as a function parameterized by $\theta\in\mathbb{R}^{m}$, where $X\in\mathbb{R}^{n}$ and $Y\in\mathbb{R}^{K}$. The $k$-th component of $\phi_{\theta}\left(x\right)$ represents the probability that sample $x$ belongs to class $k$. The objective is to adjust $\theta$ such that the model effectively “forgets” the influence of $C_{f}$ while retaining accurate predictions on $C_{r}$.
2.2. Fisher Information Matrix
The FIM is a useful tool in statistics for assessing the accuracy of parameter estimates (Kirkpatrick et al., 2017; Golatkar et al., 2020). It quantifies the impact of parameter changes on the probability distribution of observed data.
Given a probability density function $p\left(x|\theta\right)$ conditioned on the model parameter $\theta$, an FIM $I\left(\theta\right)$ is the expected value of the second derivative of the negative log-likelihood function $\ell\left(\theta\right)$, as given by
(1) $I\left(\theta\right)=\mathbb{E}\left[-\frac{\partial^{2}\ell\left(\theta\right)}{\partial\theta^{2}}\right],\ \text{where}\ \ell\left(\theta\right)=\ln p\left(x|\theta\right).$
To simplify computation, in practice, the diagonal of the FIM can approximate the second derivative of the loss function (Steven, 1993; Aich, 2021). An equivalent form of $I\left(\theta\right)$ is given by
(2) $I(\theta)=\mathbb{E}\left[\left(\left(\frac{\partial\ell\left(\theta\right)}{\partial\theta}\right)\left(\frac{\partial\ell\left(\theta\right)}{\partial\theta}\right)^{\top}\right)\right]\ .$
In this paper, the FIM is employed to quantify the importance of data to the model parameters. By comparing the FIMs of the training dataset and the to-be-unlearned dataset, we can identify the critical parameters related to the knowledge that needs to be unlearned.
3. Capturing Model-Inheritance: UMIG
In this section, we explore the topological structure of four prevalent frameworks, FL, DDPL, IL, and TL, and analyze the impact of unlearning on the inherited models within these frameworks. We propose the UMIG, which can describe the model inheritance and update relationships in these learning frameworks in a unified manner.
(a) FL vanilla and DDPL
(b) Multi-layer/server FL
(c) IL and TL
(d) Decentralized FL
Figure 1. These diagrams demonstrate various learning frameworks that can be constituted to the proposed UMIG.
3.1. Diverse Learning Frameworks and Tasks
3.1.1. Federated Learning
A central server creates a global model, which participants download to locally update the model parameters based on their local data. The participants then send back their updated parameters. The central server aggregates the parameters to update and redistribute the global model. This process repeats until the model converges (McMahan et al., 2017; Li et al., 2020b; Krauß et al., 2024). As shown in Figure 1(a), the topology of FL consists of nodes representing model states and edges representing parameter transmission and aggregation. Updates to model $A$ trigger updates in the global model $D$ and then to $E$, $F$, $G$, and $H$. FL also has other structures.
Multi-layer FL. This variant introduces an intermediate aggregation layer to traditional FL (Chu et al., 2024; Li et al., 2021), as depicted in Figure 1(b). This approach adds multiple servers that perform intermediate aggregation between the clients and the global server, thus reducing the burden on the global server, improving system scalability, and enhancing model training.
DAG-FL. As shown in Figure 1(d), multiple clients update and aggregate models hierarchically in DAG-FL (Cao et al., 2021a, b; Yu et al., 2023). Each node can receive model updates from multiple parent nodes and send the updated results to multiple child nodes. This structure enables efficient model training and more flexible communication, better adapting to complex network environments and different application needs.
3.1.2. Distributed Data-Parallel Learning
Training tasks are typically segmented into multiple sub-tasks, distributed among various computing nodes or servers for execution, as documented in (Ge et al., 2023; Dean et al., 2012; Abadi et al., 2016; Paszke et al., 2019; Li et al., 2014). Each node independently processes its designated sub-task and periodically synchronizes its model parameters with others to collectively refine the global model. This iterative process is crucial, particularly in DDPL topologies — similar to the FL topology depicted in Figure 1(a) — which require the sub-tasks $E$, $F$, and $H$, as well as the aggregated model $I$, to be updated continually with each new iteration of the global model $D$.
3.1.3. Incremental Learning
Incremental (or “Continual”) learning entails training a model initially on a dataset and then progressively updating it as new data arrives, ensuring adaptability to the latest information (Han et al., 2023; Rajapaksha et al., 2023). IL can be visualized as shown in Figure 1(c), where nodes represent different updated states of the model, and edges depict the transition between model states after data updates. If the initial state model $G$ continues to update, then the subsequent models $A$ and $B$ that receive new data, as well as their related models, also need updates.
3.1.4. Transfer Learning
TL involves pre-training a base model on one dataset and then applying this model to a new task. Typically, this process includes fine-tuning the parameters of the last few layers to better suit the new task (Sendner et al., 2023; Rajapaksha et al., 2023). The model continues to train on the specific dataset of the new task to adapt fully to its requirements. Through ongoing fine-tuning and training, the model is optimized for the best performance on the new task. The TL topology, similar to the IL topology shown in Figure 1(c), also follows a comparable model updating process.
In these diverse learning frameworks, the inheritance relationship between models is pivotal. When a base model is updated or optimized, all corresponding child models derived from it must also be updated. To accurately capture this inheritance relationship, we employ the DAG topology as an effective abstraction method to capture the dependencies among models in these learning frameworks. The use of the DAG ensures that when a single model is updated, the entire subgraph—originating from this model and encompassing all related child models—is appropriately updated.
3.2. Unified Model Inheritance Graph (UMIG)
Figure 2. The darker blue nodes are all model nodes inheriting from the starting nodes $n_{s}$ and $n_{f}$ in lighter blue. A model node $n_{\ast}$ corresponds to a model $w_{\ast}$.
To uniformly handle and analyze the inheritance and update relationships of models across various learning frameworks, including FL, DDPL, IL, and TL, we propose a DAG-based UMIG. It provides a comprehensive framework to represent components such as models, data, and tasks and their interdependencies within these frameworks. It also facilitates a structured and intuitive understanding of these complex model networks. Figure 2 illustrates the ability of UMIG to depict the inheritance relationships between models.
As illustrated in Figure 2, we define a weighted directional graph $\mathcal{G}=(N,E)$. Herein, $N=\left\{n_{1},n_{2},\ldots,n_{m}\right\}$ represents the model nodes in the network, more specifically, a set of model parameters in a specific state or task. Moreover, $E=\left\{e_{01},e_{02},\ldots,e_{mm}\right\}$ represents the inheritance relationships between model nodes, encompassing parameter transfer and task inheritance. For instance, in FL and DDPL, the edges may represent the transfer of parameters from local to global models. By contrast, in TL, the edges signify the inheritance of tasks from a pre-trained model to a fine-tuned model.
Based on the source and state of the model, we further categorize the nodes into two types:
  • Discovery Nodes refer to models trained on data with an unlearned label set, enabling these nodes to adapt to the updated classification ability within the corresponding subgraph of the DAG. The root model nodes $n_{s}$ and $n_{f}$ in Figure 2 belong to this type.
  • Inherited Nodes comprise the remaining nodes in the subgraph, except the discovery nodes. Inherited nodes point to the discovery or other inherited nodes, and focus on refining classification performance. The model nodes $n_{h},n_{j},n_{l},n_{p},n_{g},n_{i},n_{n}$ belong to this type.
3.3. UMIG-Assisted Unlearning
With an emphasis on knowledge inheritance and unlearning tasks, UMIG allows a model to retain and utilize previously learned useful information when receiving new data, achieving an effective combination of old and new knowledge. UMIG also allows the model to remove specific data for continuous optimization and performance improvement.
The UMIG-assisted unlearning consists primarily of the following two processes:
  • Locating unlearning subgraphs. Based on the data that needs to be unlearned, a breadth-first search is conducted to identify the discovery nodes, which then spawn the subgraphs associated with the discovery nodes.
  • Updating models within unlearning subgraphs. All model nodes, including discovery and inherited nodes within the identified unlearning subgraphs, are updated to unlearn the parts of their models affected by the data.
By performing a breadth-first search of the entire model network, one or more discovery nodes are identified based on unlearning requests related to specific data. These discovery nodes serve as the roots for the unlearning subgraphs. As depicted in Figure 2, the unlearning subgraphs may overlap since inherited nodes may inherit from multiple discovery nodes. For example, the discovery nodes, $n_{s}$ and $n_{f}$, lead to two overlapping subgraphs with inherited nodes. Overlapping inherited nodes, such as $n_{h},n_{j},n_{l},n_{p}$, would require the elimination of knowledge from multiple discovery nodes when unlearning due to their connections to different unlearning subgraphs. This situation is known as a multi-root scenario. In contrast, inherited nodes within one unlearning subgraph, such as $n_{g},n_{i},n_{n}$ in a single-root scenario, only need to eliminate the impact from a single discovery node.
Inheritance is not just a simple mapping from the discovery nodes to the inherited nodes. Instead, it is achieved through multiple paths and varying depths. Each path represents a route of knowledge transfer via iteration, and the number of paths and differences in depth determine the complexity of the inheritance relationship.
A complex structure emerges when multiple paths are at varying depths from an inherited node to its corresponding discovery node(s), e.g., $n_{f},n_{g},n_{i}$ or $n_{s},n_{f},n_{h},n_{j}$. If unlearning operations are conducted simultaneously in terms of depth (in other words, nodes at the same depth from the root(s) are processed simultaneously and rely on sequential dependencies), nodes such as $n_{i}$ and $n_{j}$ may require multiple updates. This situation is termed an imbalanced-path scenario, contrasting with a balanced-path scenario (e.g., $n_{f},n_{h},n_{i},n_{l}$), where each node requires only a single update due to uniform path lengths and synchronous processing at each depth. In cases where unlearning is conducted asynchronously, regardless of depth, both scenarios may require multiple updates.
Answer to RQ1: We propose the UMIG to effectively visualize model interdependence and capture the model inheritance in a structured way, leveraging the DAG structure. Assisted by UMIG, the models inheriting the knowledge to be removed can be rapidly identified. Moreover, the UMIG offers a systematic way to collect and prepare inheritance information, thus facilitating the unlearning process of inherited models. We also examine various learning frameworks of inherited models based on the numbers of their upstream models and the paths of knowledge inheritance to elucidate the emerging challenges and considerations of the unlearning process in inherited nodes due to model inheritance.
Figure 3. The unlearning process in $\mathcal{G}$ consists of five steps: 1) Preparation. The discovery nodes $\{n^{\scriptscriptstyle D}_{i}\}$ are identified by performing a breadth-first search, with unlearning data $\hat{D}^{\scriptscriptstyle D}_{i}$ containing one or more labels. Subsequently, the unlearning graph $\mathcal{\hat{G}}$ is located which root at $\{n^{\scriptscriptstyle D}_{i}\}$. 2) Calculate unlearning the unlearning FIMs $\{\hat{F}_{i}^{\scriptscriptstyle D}\}$ of $\{n^{\scriptscriptstyle D}_{i}\}$. 3) Calculate the model FIM $F_{j}$ of node $n_{j}$ in $\mathcal{\hat{G}}$. 4) Merge unlearning FIMs $\hat{F}^{\scriptscriptstyle D}_{j,k}$ for $n_{j}$ in the case of multiple discovery nodes according to the topology. 5) Update $n_{j}$ by comparing the merged unlearning FIM $\hat{F}_{j}^{\scriptscriptstyle M}$ and $F_{j}$. Steps 2, 3, 4 and 5 apply to all nodes in $\mathcal{\hat{G}}$
4. FIUn-based Parallel Unlearning
This section introduces a novel parallel unlearning method, Fisher Inheritance Unlearning (FIUn), designed to address the challenges associated with machine unlearning in large-scale UMIG model networks. A trusted learning network is considered where all learning contributors follow the designed unlearning process. The FIUn method operates under the general expression provided by the UMIG and leverages FIMs. It initiates the unlearning process by identifying specific model parameters that need updates, making the process more systematic. Derived from models in the unlearning graph, the training data, and the requested unlearning data, the FIMs are crucial in pinpointing essential model parameters and recommending their updated values.
Notably, the FIUn method allows for the independent application of unlearning steps to each impacted model, facilitating a swift and effective parallel unlearning across extensive model networks. The FIUn method is depicted in Figure 3, and the procedural details are given in Algorithm 1.
1
2Input: Model network $\mathcal{G}=(N,E)$, Unlearing task $\mathcal{T}$
3Parameter: $\gamma$, $\tau$, $\eta$
4Use breadth-first search to identify discovery nodes $\{n^{\scriptscriptstyle D}_{i}\}$ and consequential unlearning graph $\mathcal{\hat{G}}$
5 for $n^{\scriptscriptstyle D}_{i}$ in $\{n^{\scriptscriptstyle D}_{i}\}$ parallel do
6       Calculate unlearning FIM $\hat{F}^{\scriptscriptstyle D}_{i}$ with Eq. (3)
7 end for
8
9for $n_{j}$ in $\mathcal{\hat{G}}$ parallel do
10       Calculate model FIM $F_{j}$ with Eq. (4)
11       Merge unlearning FIMs $\{\hat{F}^{\scriptscriptstyle D}_{j,k}\}$ with Eq. (5)
12       Update model parameters of node $n_{j}$ with Eq. (6)
13 end for
14
Algorithm 1 Fisher Inheritance Unlearning (FIUn)
4.1. Preparation
Upon receiving an unlearning task $\mathcal{T}$, accompanied by the unlearning dataset $\hat{D}$, FIUn initiates by conducting a breadth-first search from the starting node on the model network $\mathcal{G}$ to identify the discovery nodes $\{n^{\scriptscriptstyle D}_{i}\}$. The discovery nodes are the first nodes exhibiting significant recognition of the unlearning data, with $n^{\scriptscriptstyle D}_{i}$ representing the $i$-th discovery node. Subsequently, FIUn delineates the unlearning graph $\mathcal{\hat{G}}$, which is rooted at the identified discovery nodes. FIUn then updates all nodes in the unlearning graph to unlearn the requested unlearning knowledge.
4.2. FIUn
The main idea of FIUn is to identify critical model parameters by comparing the FIM from the unlearning dataset with the FIM from the training dataset. The differences between these matrices highlight the necessary adjustments in model parameters to achieve effective unlearning. The FIUn process unfolds as follows.
4.2.1. Calculate unlearning FIMs of discovery nodes
The FIUn method initiates by assessing the impact of the unlearning data on the model parameters of the discovery nodes, utilizing the FIM as specified in Eq. (3). This assessment is conducted using the model parameters and the unlearning data (see line 5, Algorithm 1).
The unlearning FIM for the $i$-th identified discovery node, $n^{\scriptscriptstyle D}_{i}$, is denoted by $\hat{F}^{\scriptscriptstyle D}_{i}$ and is calculated as
(3) $\hat{F}^{\scriptscriptstyle D}_{i}\!\!=\!\mathbb{E}\!\left[\!\left(\!\!\left(\frac{\partial\ln p(\hat{D}_{i}^{\scriptscriptstyle D}|w^{\scriptscriptstyle L}_{i})}{\partial w^{\scriptscriptstyle L}_{i}}\right)\!\!\!\left(\frac{\partial\ln p(\hat{D}_{i}^{\scriptscriptstyle D}|w^{\scriptscriptstyle L}_{i})}{\partial w^{\scriptscriptstyle L}_{i}}\!\right)^{\!\!\!\intercal}\right)\!\!\Bigg{|}_{w^{\ast}_{\hat{D}_{i}^{\scriptscriptstyle D}}}\!\right],$
where $\hat{D}^{\scriptscriptstyle D}_{i}$ represents the unlearning data for $n^{\scriptscriptstyle D}_{i}$, $w^{\scriptscriptstyle L}_{i}$ denotes the last layer parameters of the model $w_{i}$ at node $n^{\scriptscriptstyle D}_{i}$, and $w^{\ast}_{\hat{D}_{i}^{\scriptscriptstyle D}}$ refers to the optimal parameters learned on $\hat{D}_{i}^{\scriptscriptstyle D}$.
The unlearning FIM highlights the importance of each parameter in the last layer with respect to the specific unlearning task (Lee et al., 2023; Zhang et al., 2021). In FIUn, the FIM calculation is restricted to the last layer parameters while other layers remain frozen, aimed at reducing the computational complexity associated with FIM calculations, which is a challenge in traditional FIM-based unlearning methods. Using the last layer for unlearning can focus on the most relevant features. Since the last layer directly affects the final decision of the model, adjusting its parameters can effectively facilitate unlearning goals while maintaining stable performance on data not subject to unlearning (Kim et al., 2024).
4.2.2. Calculate model FIMs of all nodes in unlearning graph
The FIUn method proceeds by calculating the model FIMs of all models within the unlearning graph to identify the impact of model parameters in each node (see line 8, Algorithm 1).
For the $j$-th node $n_{j}$ in the unlearning graph $\mathcal{\hat{G}}$, either a discovery node or an inherited node, the model FIM, denoted by $F_{j}$, is given by
(4) $F_{j}\!=\!\mathbb{E}\!\left[\!\left(\!\!\left(\frac{\partial\ln p(D_{j}|w^{\scriptscriptstyle L}_{j})}{\partial w^{\scriptscriptstyle L}_{j}}\right)\!\!\!\left(\frac{\partial\ln p(D_{j}|w^{\scriptscriptstyle L}_{j})}{\partial w^{\scriptscriptstyle L}_{j}}\right)^{\!\!\!\intercal}\right)\!\Bigg{|}_{w^{\ast}_{D_{j}}}\right]\!,$
where $D_{j}$ is the training data for $n_{j}$, $w^{\scriptscriptstyle L}_{j}$ denotes the last layer parameters of the model $w_{j}$ at node $n_{j}$, and $w^{\ast}_{D_{j}}$ denotes the optimal parameters learned on $D_{j}$.
The model FIM quantifies the knowledge about the training dataset embedded in the model parameters and highlights the significance of each parameter. Similar to the unlearning FIM, only the FIM for the last layer parameter is calculated, aiming to minimize computational complexity.
4.2.3. Merge unlearning FIMs for all nodes in the unlearning graph
The FIUn method estimates the unlearning FIM for all nodes within the unlearning graph, including both discovery and inherited nodes (see line 9, Algorithm 1). The unlearning FIMs from discovery nodes are merged to form a collective unlearning FIM for each node in the unlearning graph. This merger is crucial as the linkages between the unlearning data and parameters are embedded within the model and can persist through model inheritance (Foster et al., 2024b), especially in multi-root scenarios where an inherited node may trace back to several discovery nodes. A novel merging FIM function is introduced, denoted by $\Phi\left(\cdot\right)$, and is defined as
(5) $\hat{F}^{\scriptscriptstyle M}_{j}=\Phi\left(\{\hat{F}^{\scriptscriptstyle D}_{j,k}\}\right),$
where $\hat{F}^{\scriptscriptstyle M}_{j}$ is the merged unlearning FIM for the $j$-th node in the unlearning graph, the set $\{\hat{F}^{\scriptscriptstyle D}_{j,k}\}$ collects the unlearning FIMs of discovery nodes within $\mathcal{\hat{G}}$ that can be traced back from node $n_{j}$, and $k$ denotes the $k$-th discovery node reachable from $n_{j}$. In this paper, the element-wise maximum is employed for the merging FIM function. Specifically, $\hat{F}^{\scriptscriptstyle M}_{j}$ assembles the largest elements among the unlearning FIMs of discovery nodes in multi-root scenarios and takes the only unlearning FIM in single-root scenarios.
Figure 4. There are three discovery nodes that need to undergo the unlearning process, namely model $w_{j}$, model $w_{s}$ and model $w_{f}$. The labels to be unlearned are label 1 and label 2. However, model $w_{j}$ and model $w_{s}$ only have label 1, while model $w_{f}$ only has label 2. Model $w_{h}$ inherits from these three discovery nodes, so we merge the unlearning FIMs of the discovery nodes and perform the unlearning of label 1 and label 2 on model $w_{h}$ together.
As shown in Figure 4, the model $w_{h}$ inherits from models $w_{j}$, $w_{s}$, and $w_{f}$ and needs to remove all requested unlearning data, though each discovery node contains only a portion of the unlearning data. The FIM of each model reflects the sensitivity of its parameters to the unlearning data, which can be interpreted as parameter importance. By taking the maximum of the corresponding elements across these unlearning FIMs, the merged FIM encompasses all unlearning tasks from the discovery nodes and effectively identifies the parameters that need updating to achieve complete unlearning.
4.2.4. Update models in unlearning graph
In the final step, the FIUn method identifies which parameters need to be updated by comparing the merged FIM with the model FIM, as described in line 10 of Algorithm 1. The outcomes of this comparison are also utilized to scale the parameters. For node $n_{j}$ in the unlearning graph, the parameters are updated as
(6) $\hat{w}^{\scriptscriptstyle L}_{j,l}=\begin{cases}\min(\tau\frac{F_{j,l}}{\hat{F}^{\scriptscriptstyle M}_{j,l}},\eta)w^{\scriptscriptstyle L}_{j,l},&\text{if }\frac{\hat{F}^{\scriptscriptstyle M}_{j,l}}{F_{j,l}}>\gamma;\\ w^{\scriptscriptstyle L}_{j,l},&\text{if}\ \frac{\hat{F}^{\scriptscriptstyle M}_{j,l}}{F_{j,l}}\leq\gamma,\end{cases}$
where $w^{\scriptscriptstyle L}_{j,l}$ and $\hat{w}^{\scriptscriptstyle L}_{j,l}$ are the $l$-th parameter in the last layer before and after the unlearning update, respectively. Similarly, $F_{j,l}$ and $\hat{F}^{\scriptscriptstyle M}_{j,l}$ are the $l$-th element of the model FIM and merged unlearning FIM of node $n_{j}$, respectively. The hyperparameters, including $\tau$, $\eta$, and $\gamma$, balance the unlearning impact on the unlearning data and the accuracy of the retained data.
During the model updating phase, only the parameters of the last layer are updated while the other layers remain frozen, consistent with our design of the unlearning FIM and model FIM, where only the FIM of the last layer is calculated. Elements within the FIM indicate the significance of each parameter relative to the dataset. The threshold $\gamma$ identifies the critical parameters that are significant for the unlearning task, and these parameters are scaled down accordingly. The hyperparameter $\eta$ ensures minimal change to achieve the desired unlearning effect. Non-critical parameters are left unchanged to maintain accuracy on data that does not require unlearning.
4.3. Discussion
From the perspective of the FIM, the impact of data on specific model parameters is consistently localized at precise positions within the parameter space (Foster et al., 2024b). The stable localization of the FIM across iterations helps precisely locate the parameter locations for unlearning at the class, client, and sample levels using data tied to specific label sets. By fine-tuning $\tau$, $\eta$, and $\gamma$ in Eq. (6) to match the sample size of the class being unlearned, the task transits flexibly between the class-level (Foster et al., 2024b) and sample-level unlearning (Golatkar et al., 2020). Specifically, when the unlearning FIM mirrors the model’s FIM with a class-level setting, it becomes a client-level unlearning task. All types of unlearning are performed irrespective of the succession of inheritance relationships. Consequently, employing FIUn disrupts the sequential dependencies traditionally required for unlearning, thus enabling fully parallel unlearning processes across any number of subgraphs.
Not only can models across different subgraphs undergo parallel unlearning, but those within the same subgraph at various depths can also be unlearned in parallel, thanks to a feature referred to as Hyper-Distance in this paper. This Hyper-Distance is achieved without compromising the integrity of the inheritance relationships. Hyper-Distance ensures that both unlearned and retained knowledge is effectively managed along the inheritance paths. This efficiency is possible because the FIUn method utilizes FIMs for unlearning, enabling a class-wise inheritance approach. Each model effectively incorporates a union of label sets from its predecessors, ensuring that the inheritance is preserved. Simultaneously, by focusing on label differences, we achieve fully parallel unlearning—making depths and paths irrelevant to the process.
Answer to RQ2: The FIUn method enables independently removing knowledge from each inherited model. This method utilizes the FIM to obtain the Hyper-Distance property which breaks sequential dependencies among models, enabling fully parallel unlearning on inherited models and significantly decreasing the execution time of unlearning tasks. To address the complexity arising from multiple upstream models, the merging FIM function is developed to aggregate unlearning FIMs from those models, facilitating the efficient, one-shot removal of inherited knowledge.
We also map the FIUn method and merging FIM function to the UMIG-assisted unlearning scenarios described in Section 3.3. This demonstrates how the proposed approach efficiently facilitates unlearning through the property of Hyper-Distance. Converting Imbalance into Balance. This scenario is particularly significant due to the fact that existing unlearning methods suffer from inevitable sequential dependencies. FIUn leverages the Hyper-Distance property to circumvent the influences of depths and paths within the inheritance topology. Consequently, imbalanced-path scenarios can be managed as effectively as balanced-path scenarios, where only one update operation is needed, regardless of whether unlearning is conducted synchronously or asynchronously, i.e., only one unlearning FIM needs to be computed for each inherited node.
Converting Multi-root into Single-root. The merging FIM function is proposed to refine multi-root scenarios. This function merges the FIMs of the corresponding unlearned label sets from multiple discovery nodes into a single FIM, as shown in Figure 4. Specifically, overlapping inherited nodes require only one update operation to eliminate the impact of multiple discovery nodes by conducting a parameter-wise merger, where only the most impactful parameters take effect in this FIM. This function eliminates the need for multiple comparisons among various unlearning FIMs in multi-root scenarios, hence enhancing unlearning efficiency.
4.4. Implementation
4.4.1. System Setting
In practice, FIUn can be handled differently depending on the system architecture, as described in the following:
  • Centralized: Model FIMs are computed and attached along with the model being published and sent to the central coordinator. When an unlearning request is initiated, the central server performs unlearning by merging the unlearning FIMs of the discovery nodes and executes unlearning in parallel.
  • Decentralized: The discovery nodes publicly share Each unlearning FIM. Owners of each inherited node can then combine these unlearning FIMs with their own local models at will, according to specific unlearning requirements. This decentralized approach enables efficient and effective unlearning.
4.4.2. Complexity
The computational complexity of an unlearning task comprises the complexities of Steps 1, 3, and 4 (i.e., calculate unlearning FIM, merge unlearning FIMs and update models). The complexity of Step 2 is not considered, as the computation of the model FIM can be performed beforehand. The overall complexity for the unlearning task is $\mathcal{O}(|\hat{D}||L|+|\{n^{\scriptscriptstyle D}_{i}\}|+|L|)$. Breaking down the complexities, Eq. (3) and Eq. (4) update $|L|$ parameters per data entry, with complexities $\mathcal{O}(|\hat{D}^{\scriptscriptstyle D}_{i}||L|)$ and $\mathcal{O}(|D_{j}||L|)$, respectively, where $|\hat{D}^{\scriptscriptstyle D}_{i}|\leq|\hat{D}|$ and $|L|$ is the number of parameters in the last layer. The term $\mathcal{O}(|D_{j}||L|)$ is omitted as model FIMs can be calculated along with the model training prior to any unlearning tasks. Eq. (5) includes a comparison and merger of $K$ unlearning FIMs, leading to a complexity of $\mathcal{O}(K)$, where $K$ is the number of traceable discovery nodes in the unlearning graph $\mathcal{\hat{G}}$ from node $n_{j}$ and $K\leq|\{n^{\scriptscriptstyle D}_{i}\}|$. Eq. (6) updates $|L|$ model parameters, leading to a complexity of $\mathcal{O}(|L|)$. As the computation of the unlearning FIM for discovery nodes and the model updates can proceed in parallel, the overall complexity of the unlearning tasks is $\mathcal{O}(|\hat{D}||L|+|\{n^{\scriptscriptstyle D}_{i}\}|+|L|)$.
5. Experimental Settings
This section first presents the datasets used in the experiments, the models selected, and the performance metrics for accuracy evaluation, followed by the topology setups and the benchmarks for unlearning to evaluate the FIUn method.
Table 1. Baseline of Federated Unlearning Performance On CIFAR-100 Unlearning Speed Comparison
Model #$C_{f}$ Cumulative Unlearning Time (s)
Gradient ascent Finetune Distill Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
AlexNet 1 0.97 1.05 1.26 0.54 0.75 1.08 82.5 153.65 294.32 29.70 59.06 89.62 0.09 0.11 0.13
2 0.99 1.17 1.39 0.57 0.77 1.15 80.49 155.63 315.52 28.69 57.81 84.97 0.11 0.16 0.13
10 1.1 1.27 1.68 0.53 1.95 2.31 75.3 147.6 326.8 21.96 42.37 63.82 0.11 0.14 0.15
ResNet18 1 2.45 2.89 3.42 0.74 1.32 1.74 70.31 129.44 247.6 30.36 61.40 90.94 0.68 1.00 0.99
2 2.02 2.2 2.78 0.76 1.22 1.79 76.55 139.55 259.53 28.30 57.39 84.06 0.69 1.03 1.09
10 2.02 2.58 3.67 0.85 1.54 1.84 75.12 150.35 267.34 22.03 44.83 66.30 0.76 1.14 1.04
(a) FL in ResNet18 of #$C_{f}=1$
(b) FL in ResNet18 of #$C_{f}=10$
(c) FL in AlexNet of #$C_{f}=1$
(d) FL in AlexNet of #$C_{f}=10$
Figure 5. Baseline of Federated Unlearning Performance On CIFAR-100.
5.1. Datasets and Models
Datasets. The experiment uses two datasets to evaluate the proposed FIUn method. These two datasets are benchmarks for image classification tasks, encompassing diverse image classification tasks.
  • CIFAR100 (Krizhevsky et al., 2009). This dataset consists of 60,000 color images in 100 classes, with 600 images per class. The images are of size $32\times 32$ pixels. The dataset is divided into 50,000 training images and 10,000 test images. Each class in the CIFAR-100 dataset belongs to one of 20 superclasses. There is a coarse label (superclass) and a fine label (specific class) per image.
  • TinyImageNet (Le and Yang, 2015). This dataset contains 200 classes, each with 500 training images, 50 validation images, and 50 test images, with a total of up to 100,000 images. The images are resized to $64\times 64$ pixels, making them smaller than those in the original ImageNet dataset but larger than those in CIFAR-100.
Models. We use the AlexNet model (Iandola et al., 2016), ResNet18 model (Odusami et al., 2021) and DenseNet161 model (Talo, 2019) to evaluate the impact of our proposed FIUn method on the accuracy of the unlearned label set $C_{f}$ and the retained label set $C_{r}$.
Performance Metrics. We use the training datasets to check the accuracy of the model to evaluate its practicality in experiments. This is a common and reasonable approach since it directly evaluates if the removed information still influences the model, as widely accepted in the field (Tarun et al., 2023; Kim et al., 2024; Golatkar et al., 2020).
  • Accuracy on unlearned labels ($AD_{f}$). The accuracy of the unlearned label set in the unlearning model ideally would be close to zero. We use the training datasets to test the effectiveness of unlearning.
  • Accuracy on retained labels ($AD_{r}$). The accuracy of a retained label set of the unlearning model. It is expected to be close to the accuracy of the original model before unlearning.
  • Cumulative unlearning time. The cumulative time required for each model to unlearn labels during the training process in machine learning and deep learning.
  • Accuracy difference index ($\bigtriangleup_{acc}$). The difference between $AD_{r}$ and $AD_{f}$ also serves as a direct metric to evaluate unlearning performance. The greater the difference, the more effective the unlearning is, with a maximum 1.
Table 2. Federated Unlearning Performance On TinyImageNet
Model #$C_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
DenseNet161 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.81 99.99 99.99 98.80 88.03 77.64 76.02 999.41 1996.43 2997.64 2.51 4.59 4.95
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.99 99.99 99.80 99.99 99.99 99.51 83.01 71.43 71.00 1004.64 2006.31 3006.14 3.16 4.69 4.89
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00
6 ${AD}_{r}\uparrow$ 99.99 99.99 99.96 99.99 99.99 99.16 80.00 67.67 66.58 1019.43 2020.16 3018.64 3.10 4.96 4.75
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.99 99.99 99.72 99.71 99.99 99.99 71.68 58.01 55.80 1031.43 2029.64 3030.64 2.95 4.85 4.89
${AD}_{f}\downarrow$ 99.99 99.99 98.82 0.00 0.00 0.00 0.00 0.00 0.00
ResNet18 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.73 99.97 99.46 112.95 216.39 314.29 1.35 3.68 3.42
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.17 99.98 99.98 109.89 209.31 310.42 1.63 3.26 3.57
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
6 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 99.54 99.98 99.98 107.39 208.12 305.32 1.46 3.37 3.51
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.98 99.99 99.99 104.21 201.36 302.74 1.52 3.21 3.36
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
Table 3. Incremental Unlearning Performance On TinyImageNet
Model #$C_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
DenseNet161 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.81 99.99 99.99 98.80 89.76 78.50 73.87 991.43 1995.36 3989.47 2.75 4.85 4.43
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.99 99.99 99.80 99.99 99.99 99.51 83.98 65.87 59.11 1006.32 2004.75 4009.41 2.64 4.43 4.64
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00
6 ${AD}_{r}\uparrow$ 99.99 99.99 99.96 99.99 99.99 99.16 86.81 61.02 55.04 1019.43 2019.50 4019.46 2.36 4.47 4.74
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.99 99.99 99.72 99.99 99.99 99.99 85.95 54.61 45.46 1035.36 2041.78 4039.75 2.65 4.46 4.75
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
ResNet18 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 98.66 96.33 96.07 120.32 221.75 439.85 1.37 3.75 3.43
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 98.84 96.34 96.70 115.32 217.33 435.17 1.41 3.34 3.64
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
6 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 98.82 99.99 98.59 110.32 211.35 431.37 1.47 3.46 3.75
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.33 97.51 98.45 108.32 207.43 428.43 1.32 3.56 3.57
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
5.2. Learning Framework Setup and Benchmarks
Figure 6. Models topologies considered in the experiments.
Targeted Frameworks. We detailed setups for four learning frameworks under the UMIG with varying dataset and label distribution conditions.
  • Federated Unlearning. This process uses the structure shown in Figure 6. The CIFAR100 and TinyImageNet datasets are each randomly divided into five parts. When the model $w_{g}$ contains labels that need to be unlearned, the models that inherit from it need to perform the unlearning operation. We select models $w_{g}$, $w_{a}$, and $w_{b}$ for experimental demonstration.
  • Distributed Data-Parallel Unlearning. This process uses the structure shown in Figure 6. The CIFAR100 and TinyImageNet datasets are each randomly divided into three parts, each executing a sub-task in parallel. When the model $w_{g}$ contains labels that need to be unlearned, each sub-task performs the unlearning operation in parallel. We select models $w_{g}$, $w_{a}$, and $w_{b}$ for experimental demonstration.
  • Incremental Unlearning. This process uses the structure shown in Figure 6. In the CIFAR100 dataset, the first model is trained on data with labels 0-90, and each inheriting model’s training dataset progressively includes two additional classes. In the TinyImageNet dataset, the first model is trained on data with labels 0-190, and similarly, each inheriting model’s training dataset progressively includes two additional classes. We select models $w_{g}$, $w_{a}$, and $w_{b}$ for experimental demonstration.
  • Transfer Unlearning. Since TL involves fine-tuning the last layer parameters, adjustments are focused on the last layer of the model. Experiments, as illustrated in Figure 6, involve the CIFAR100 dataset where a model $w_{g}$ is trained on labels 0-90 and then transferred to models $w_{a}$ and $w_{b}$ trained on labels 0-92 and 0-98, respectively. Similarly, for the TinyImageNet dataset, model $w_{g}$ is trained on labels 0-190 and then transferred to models $w_{a}$ and $w_{b}$, which are trained on labels 0-192 and 0-198, respectively.
Benchmark. The experiment specifically focuses on class-level unlearning tasks for clarity of presentation. We re-train the models that include unlearned labels and subsequent models influenced by these labels due to inheritance. We remove the data with unlearned labels from each model’s dataset. The remaining data is then used to individually train each model. The models are trained sequentially according to their inheritance relationships.
Hyperparameters. The hyperparameters used in the experiment are shown in Appendix .1. The table lists the learning rates, optimizers, and other hyperparameters for different models, as well as the hyperparameters of our proposed FIUn method. Throughout the experimental process, all the hyperparameters remain unchanged to ensure a fair comparison of our method among different models.
Table 4. Transfer Unlearning Performance On CIFAR-100
Model #$C_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
AlexNet 1 ${AD}_{r}\uparrow$ 97.16 95.61 92.20 98.17 95.98 91.32 91.40 86.24 83.99 20.31 40.49 42.36 0.09 0.11 0.11
${AD}_{f}\downarrow$ 96.66 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
2 ${AD}_{r}\uparrow$ 97.16 95.61 92.20 96.88 96.68 93.16 82.36 80.47 79.99 21.20 40.56 41.92 0.08 0.13 0.11
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 97.16 95.61 92.20 94.67 96.45 93.45 71.59 71.43 76.06 20.30 39.76 40.60 0.09 0.11 0.11
${AD}_{f}\downarrow$ 96.61 98.30 98.30 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 97.16 95.61 92.20 99.99 99.99 99.99 75.97 69.56 76.15 19.42 39.46 39.35 0.08 0.13 0.11
${AD}_{f}\downarrow$ 98.57 99.28 99.28 0.00 0.00 0.00 20.28 20.28 20.28
ResNet18 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 89.09 93.51 95.22 22.21 41.52 42.93 0.15 0.24 0.24
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
2 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 85.89 94.89 95.79 21.19 41.72 43.20 0.17 0.23 0.26
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.96 99.99 99.99 99.99 99.99 99.99 80.20 94.82 97.03 22.03 42.26 43.25 0.17 0.25 0.25
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.96 99.99 99.99 99.99 99.99 99.99 75.36 98.07 98.96 19.10 42.60 41.36 0.16 0.27 0.27
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
Table 5. Distributed Data-Parallel Unlearning Performance On CIFAR-100
Model #$C_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
AlexNet 1 ${AD}_{r}\uparrow$ 98.01 99.96 99.96 99.96 99.96 99.96 97.23 98.85 97.66 30.13 31.26 31.14 0.09 0.14 0.23
${AD}_{f}\downarrow$ 99.99 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00
2 ${AD}_{r}\uparrow$ 99.96 99.96 99.96 99.96 99.96 99.96 96.14 97.07 96.71 26.43 26.47 26.41 0.11 0.16 0.27
${AD}_{f}\downarrow$ 99.96 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 98.01 99.96 99.96 99.96 99.96 99.96 94.58 90.41 90.97 24.33 24.36 24.37 0.10 0.14 0.23
${AD}_{f}\downarrow$ 99.09 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 98.01 99.96 99.96 99.96 99.96 99.96 91.53 76.19 80.70 21.03 22.96 22.10 0.10 0.14 0.24
${AD}_{f}\downarrow$ 97.15 99.96 99.96 0.00 0.00 0.00 9.73 0.00 11.08
ResNet18 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 91.87 94.67 93.67 32.53 31.95 32.18 0.30 0.34 0.64
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
2 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 89.94 94.58 92.40 28.28 27.22 28.47 0.30 0.34 0.66
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.96 89.20 88.52 99.96 99.96 86.44 93.95 89.26 91.77 25.37 24.97 25.13 0.31 0.35 0.66
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.96 89.20 88.52 99.96 99.96 99.96 90.44 95.79 94.19 22.60 21.92 24.19 0.34 0.34 0.68
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 9.73 0.00 0.00
Table 6. Federated Unlearning Performance On TinyImageNet With Multiple Label Distribution
Model #flimit-$\#\cap_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
DenseNet161 60% $AD_{r}\uparrow$ 99.99 99.99 99.81 98.80 99.99 99.99 96.39 84.75 77.89 3000.66 1998.46 994.32 4.69 2.39 2.30
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00
40% $AD_{r}\uparrow$ 99.50 99.99 99.80 99.51 99.99 99.99 96.64 85.40 85.00 3009.53 1010.45 2011.14 5.12 2.69 2.14
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00
20% $AD_{r}\uparrow$ 99.99 99.99 99.96 99.16 99.99 99.99 96.22 81.49 78.83 3022.64 2020.67 1021.34 4.98 2.64 2.59
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00
0% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.71 96.01 83.04 79.15 3039.64 2041.32 1046.14 4.96 2.14 2.50
${AD}_{f}\downarrow$ 99.99 99.99 98.82 0.00 0.00 0.00 0.00 0.00 0.00
ResNet18 60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 95.23 98.54 98.44 349.32 246.43 123.74 3.16 1.65 1.68
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
40% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 96.98 99.44 99.49 355.32 247.64 128.43 2.97 1.46 1.59
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 94.43 98.86 99.38 354.83 246.43 129.35 2.89 1.54 1.38
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
0% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 97.36 99.34 99.32 351.38 244.34 125.32 3.10 1.69 1.39
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
Table 7. Distributed Data-Parallel Unlearning Performance On CIFAR-100 With Multiple Label Distribution
Model #flimit-$\#\cap_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
DenseNet161 60% $AD_{r}\uparrow$ 99.99 99.99 99.81 98.80 99.99 99.99 95.18 97.07 99.99 999.41 996.43 997.64 0.15 0.08 0.09
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00
40% $AD_{r}\uparrow$ 99.99 99.99 99.80 99.51 99.99 99.99 90.53 97.07 98.82 1004.64 1006.31 1006.14 0.19 0.09 0.11
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00
20% $AD_{r}\uparrow$ 99.99 99.99 99.96 99.16 99.99 99.99 95.29 97.07 99.89 1019.43 1020.16 1018.64 0.18 0.07 0.10
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00
0% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 87.15 97.07 96.36 1031.43 1029.64 1030.64 0.18 0.10 0.11
${AD}_{f}\downarrow$ 99.99 99.99 98.82 0.00 0.00 0.00 0.00 0.00 0.00
ResNet18 60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 97.62 88.94 87.87 30.36 33.18 37.31 0.60 0.32 0.29
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
40% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 96.50 89.35 87.89 28.30 29.13 52.33 0.59 0.34 0.30
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 95.31 88.17 93.06 26.39 23.37 56.61 0.58 0.28 0.30
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
0% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 88.87 88.86 89.16 22.03 24.36 46.42 0.64 0.37 0.32
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
5.3. Label Category Settings
The unlearned label categories are represented as #$\#C_{f}$. The distribution of labels among different models may also vary in different data distributions. For instance, model $w_{a}$ aims to unlearn labels 1, 2, 3, and 4, while model $w_{b}$ intends to unlearn labels 2, 3, 5, and 6. In this case, there is an overlap between the unlearned label categories of model $w_{a}$ and model $w_{b}$, specifically labels 2 and 3. We refer to the degree of overlapping among these unlearned label categories as #flimit-$\#\cap_{f}$.
The unlearning process is applied to individual categories and various combinations of labels. For the CIFAR-100 dataset, the combinations we select include 1, 2, 4, and 10 categories. For the TinyImageNet dataset, the selected combinations include 1, 4, 6, and 10 categories. The overlap of the unlearned label categories is divided into different proportions, including 60%, 40%, 20%, and 0%.
6. Evaluation
In this section, we use NVIDIA 4090 GPU to set up a test platform that demonstrates the impact of the FIUn method on the accuracy of unlearned and retained labels in four frameworks: FL, DDPL, IL and TL. When assessing accuracy and time consumption across these learning frameworks during the model unlearning process, we compare our proposed FIUn method with existing unlearning methods and use re-training as a benchmark. We also analyze the impact of the inheritance depth and parameter layers of the models on unlearning effectiveness.
(a) FL in CIFAR100 ResNet18
(b) FL CIFAR100 by AlexNet
(c) FL TinyImageNet by ResNet18
(d) FL TinyImageNet by DenseNet161
Figure 7. Overlap #$\cap_{f}$ analysis, #$C_{f}$=15.
6.1. General Label Unlearning Analysis
We consider existing emerging unlearning methods for comparison: 1) Fine-tuning, which adjusts the original model using the remaining dataset (Odusami et al., 2021), 2) Gradient ascent (GA), employing negative gradients for unlearning (Zhang et al., 2024), and 3) Distill, which distills knowledge from the original model into a student model using the remaining data (Sinha et al., 2023).
Performance comparisons are depicted in Figure 5 for unlearning accuracy $\bigtriangleup_{acc}$ and detailed in Table 1 for unlearning speed. The results demonstrate that our method not only enhances unlearning speed by up to 99% over alternative unlearning methods but also consistently outperforms them in unlearning accuracy.
These results indicate that our method exhibits outstanding effectiveness and efficiency in handling the complexity of model unlearning. In what follows, we evaluate the proposed parallel unlearning method in comparison with the most robust method—re-training. Within all considered learning frameworks, our proposed FIUn method achieves a maximum increase in unlearning speed of 99% compared to the re-training benchmark across different datasets and models. The detailed results are summarized in Tables 2 to 5, with more experiments provided in Appendix .2.1.
1) Single-label unlearning. For all frameworks and model types in CIFAR100, experimental results show that the $AD_{f}$ metric reaches 0, strongly proving that our method effectively unlearns single-class labels. Meanwhile, the average $AD_{r}$ metric is as high as 94.53%, further confirming our method’s efficiency in retaining labels. For all frameworks and model types in TinyImageNet, the $AD_{f}$ metric also reaches 0, and the average $AD_{r}$ metric reaches 79.49%. This significantly validates the effectiveness of our method.
2) Multi-label unlearning. For the CIFAR100 dataset, with #$C_{f}=2,4,10$, the average $AD_{f}$ metric across all frameworks and model types is 1.93%, while the average $AD_{r}$ metric reaches 86.01%. For the TinyImageNet dataset, the average $AD_{f}$ metric is 0.21%, and the average $AD_{r}$ metric is 83.54%. These results further demonstrate the effectiveness of our method. On the other hand, for both the IL and TL frameworks, as the number of unlearning label classes increases, the $AD_{f}$ metric for all models approaches 0. However, the $AD_{r}$ performance of the AlexNet and DenseNet161 models shows a downward trend, while the $AD_{r}$ performance of the ResNet18 model remains stable.
Further observation reveals that the number of parameters in the last layer of DenseNet161 is 4.3 times that of ResNet18, and the number of parameters in the last layer of AlexNet is 7.9 times that of ResNet18. These results indicate that as the number of model parameters and the classes of forgotten labels increases, the $AD_{r}$ performance may decline. This further suggests that while our proposed method effectively unlearns specified labels, it can somewhat affect the accuracy of retained labels.
(a) FL TinyImageNet by ResNet18
(b) FL CIFAR100 by ResNet18
(c) FL TinyImageNet by DenseNet161
(d) FL CIFAR100 by DenseNet161
(e) IL TinyImageNet by ResNet18
(f) IL CIFAR100 by ResNet18
(g) IL TinyImageNet by DenseNet161
(h) IL CIFAR100 by DenseNet161
Figure 8. Inheritance depth analysis #$C_{f}=4$.
(a) FL AlexNet in CIFAR100
(b) FL CIFAR100 by ResNet18
(c) IL CIFAR100 by ResNet18
(d) IL CIFAR100 by AlexNet
(e) FL TinyImageNet by DenseNet161
(f) FL TinyImageNet by ResNet18
(g) IL TinyImageNet by ResNet
(h) IL TinyImageNet by DenseNet161
Figure 9. Unlearning layer analysis #$C_{f}$=15.
6.2. Merged Label Unlearning Analysis
This section assesses the applicability of the merging FIM function to models with various distributions of unlearned labels. We analyze the unlearning performance of different types of models across various datasets within the FL and DDPL frameworks. As demonstrated in Tables 6 and 7, with more experiments provided in Appendix .2.2, the average $AD_{r}$ is 89.72%, and the average $AD_{f}$ is 5.62%, confirming the unlearning effectiveness of our method. We further validate that in the FIM, the impact of data on model parameters remains relatively stable.
We also analyze whether our proposed merging FIM function can handle the merged label unlearning task within the FL framework. As illustrated in Figure 7, in the AlexNet and DenseNet161 models, our method shows an average performance improvement of 88.54%, compared to using FIM directly. In the ResNet18 model, our method’s performance closely matches that of FIM. Additionally, our proposed merged label unlearning is, on average, 2.3 times faster than general label unlearning. This further validates the effiency of our method.
6.3. Inherited Depth of Model Analysis
We further examine whether the performance of FIUn is influenced by the increasing depth of inheritance within the FL and IL frameworks. All experiments are conducted with the same hyperparameters.
  • Federated Learning. In the FL framework, the CIFAR100 and TinyImageNet datasets are randomly divided into five segments. Each model trains using one segment of the data, adopting a binary tree structure as depicted in Figure 6. We evaluate whether our method can effectively unlearn specified labels as the depth of the binary tree increases.
  • Incremental Learning. In the IL framework, as demonstrated in Figure 6, the starting model $w_{g}$ contains data from 170 classes. With each incremental learning step, one class of label data is added to assess whether our method can effectively unlearn specified labels.
As shown in Figure 8 (i.e., single-class unlearning, see Appendix .4), we observe that under both frameworks, various datasets, and models with different unlearned labels, the performance of the proposed algorithm remains consistent regardless of the increased inherited depth. This consistency across both frameworks and models supports the notion that the impact of data on model parameters remains stable at fixed positions. Such stability ensures that our method maintains high performance even as the depth increases.
6.4. Analysis of Parameter Layer Selection in Models
We analyze how the number of model layers used to calculate the FIM impacts the performance of our algorithm. We conduct experiments within both FL and IL frameworks, with 15 label categories selected as unlearned labels. The DAG diagrams for these frameworks are illustrated in Figures 6 and 6. We calculate the FIM using the last layer, the last four layers, the last seven layers, and all layers to assess unlearning performance.
As depicted in Figure 9, we observe that an increase in the number of layers used leads to diminishing unlearning performance across various datasets and models. This decline is attributable to the increase in the layer parameters, which inversely affects the $AD_{r}$ and $AD_{f}$ metrics. Moreover, a larger number of layers prolongs the time needed to compute the FIM. This finding underscores the effectiveness of our method, which relies on calculations from only the last layer.
7. Related Work
Although FL (Li et al., 2020b), DDPL (Li et al., 2020c), IL (Van de Ven et al., 2022), and TL (Weiss et al., 2016) have each made significant strides in addressing specific challenges, such as data privacy protection (Li et al., 2021), model training efficiency (Dean et al., 2012), knowledge retention and unlearning (Han et al., 2023), and cross-domain knowledge transfer (Rajapaksha et al., 2023), systematic exploration that combines these learning frameworks remains sparse. Most studies have merely attempted simple combinations of a few methods (Dai and Meng, 2023; Yoon et al., 2021; Dai et al., 2019; Miao et al., 2023) without touching the core of their commonality–model inheritance. This characteristic plays a crucial role in environments where unlearning is conducted on a large-scale. As data and models grow, the computational overhead increases exponentially with the depth of inheritance, while time consumption grows linearly with the depth, even when unlearning can be executed simultaneously at the same depth starting from the unlearning origin. This holds true regardless of the existing unlearning methods utilized, which are respectively discussed in the following.
Federated Unlearning. FedEraser (Liu et al., 2020) used historical parameter updates from the central server to reconstruct the deleted model. Wu et al. (Wu et al., 2022a) subtracted historical updates and used knowledge distillation to restore model performance without client data. FRU (Yuan et al., 2023) recalled and calibrated historical updates to eliminate user contributions and accelerate federated recommendation reconstruction. These methods primarily relied on recalling and calibrating historical parameter updates of the model to achieve deletion or re-training. Halimi et al. (Halimi et al., 2022) reversed the learning process by maximizing local empirical loss and used PGD to perform deletion on the client side, while Wu et al. (Wu et al., 2022b) used class, client, and sample decoupling learning combined with Stochastic Gradient Ascent (SGA) and Elastic Weight Consolidation (EWC) to achieve deletion. These methods achieved deletion by optimizing the model’s loss function or gradients.
Liu et al. (Liu et al., 2022) utilized first-order Taylor expansion and a diagonal empirical Fisher information matrix for accelerated re-training. Wang et al. (Wang et al., 2022) applied CNN channel pruning to remove specific category information, while FFMU (Che et al., 2023) employed nonlinear function analysis to enhance model deletion processes. These approaches streamline deletion or re-training by altering the model structure or employing targeted techniques. Additionally, KNOT (Su and Li, 2023) optimized re-training by introducing clustering aggregation and framing the client clustering issue as a dictionary minimization problem.
Distributed Data-Parallel Unlearning. In DDPL knowledge unlearning, handling the unlearning process is restricted to the data shard on the specific device, followed by model aggregation and iterative training. If unlearning is required from a specific shard, only the model trained on that shard is re-trained. Approximate machine unlearning directly modifies trained model parameters, employing methods such as 1) Gradient-based approaches (Fan et al., 2023; Golatkar et al., 2020; Neel et al., 2021), which update model gradients to gradually diminish the data’s influence, and 2) Data perturbation (Tarun et al., 2023; Goel et al., 2024), which introduces noise to reduce specific data points’ impact on the model. These methods facilitate effective knowledge unlearning in DDPL.
Incremental Unlearning. Research on unlearning on IL is still nascent, yet the Projected-Gradient Unlearning method (PGU) (Hoang et al., 2024) shows promise. PGU computes the gradients for the unlearning dataset and projects these onto the Core Gradient Space (CGS) derived from the training set. It then subtracts this projected component from the original gradients, positioning the residuals in a space orthogonal to the CGS, and updates the model parameters with these adjusted gradients. Effective for one-time data deletions and batch-wise unlearning, PGU is well-suited for ongoing incremental learning processes.
Transfer Unlearning. Research on unlearning specific target task knowledge has been relatively scarce in TL. To address this challenge, (Sepahvand et al., 2024) proposed fine-tuning pre-trained models through data selection. Subsequently, the entire network was fine-tuned on the target task using gradient descent.
What Sets Our Solution Apart. The above-mentioned existing studies on unlearning methods typically endure high computational complexity, significant resource demands, and implementation difficulties. These methods are time-consuming and resource-intensive, especially when deployed in large-scale environments, as they require accurate tracking and sequential updating of numerous model parameters. By contrast, the proposed FIUn method fully parallelizes unlearning, significantly enhancing computational efficiency, resource utilization, and adaptability. It effectively addresses the challenges faced by existing unlearning methods in inheritance-oriented model networks.
8. Conclusion
In this paper, we developed a new parallel unlearning framework for models with dependence and inheritance. To do this, we introduced the UMIG, which can effectively abstract model inheritance, e.g., in FL, DDPL, IL, and TL, using a DAG topology. We also developed the FIUn method and the merging FIM function, which can quickly locate impacted models and conduct parallel unlearning through the UMIG. Experimental results on various models and datasets showed that our method achieves complete unlearning for single-class labels, while maintaining an average accuracy of 94.53% for retained labels. For multi-class labels, the unlearning accuracy is only 1.07%, with retained label accuracy at 84.77%. Our method is 99% faster than existing unlearning methods, efficiently handling model updates regardless of inheritance depth.

### v4

@@
-Fishers Harvest Parallel Unlearning in Inherited Model Networks
-Abstract.
-Unlearning in various learning frameworks remains challenging, with the continuous growth and updates of models exhibiting complex inheritance relationships.
-This paper presents a novel unlearning framework, which enables fully parallel unlearning among models exhibiting inheritance.
-A key enabler is the new Unified Model Inheritance Graph (UMIG), which captures the inheritance using a Directed Acyclic Graph (DAG).
-Central to our framework is the new Fisher Inheritance Unlearning (FIUn) algorithm, which utilizes the Fisher Information Matrix (FIM) from initial unlearning models to pinpoint impacted parameters in inherited models.
-By employing FIM, the FIUn method breaks the sequential dependencies among the models, facilitating simultaneous unlearning and reducing computational overhead.
-We further design to merge disparate FIMs into a single matrix, synchronizing updates across inherited models.
+Parallel Unlearning in Inherited Model Networks
+Abstract
+Unlearning is challenging in generic learning frameworks with the continuous growth and updates of models exhibiting complex inheritance relationships.
+This paper presents a novel unlearning framework that enables fully parallel unlearning among models exhibiting inheritance.
+We use a chronologically Directed Acyclic Graph (DAG) to capture various unlearning scenarios occurring in model inheritance networks.
+Central to our framework is the Fisher Inheritance Unlearning (FIUn) method, designed to enable efficient parallel unlearning within the DAG.
+FIUn utilizes the Fisher Information Matrix (FIM) to assess the significance of model parameters for unlearning tasks and adjusts them accordingly.
+To handle multiple unlearning requests simultaneously, we propose the Merging-FIM (MFIM) function, which consolidates FIMs from multiple upstream models into a unified matrix.
+This design supports all unlearning scenarios captured by the DAG, enabling one-shot removal of inherited knowledge while significantly reducing computational overhead.
 Experiments confirm the effectiveness of our unlearning framework.
-For single-class tasks, it achieves complete unlearning with 0% accuracy for unlearned labels while maintaining 94.53% accuracy for retained labels on average.
-For multi-class tasks, the accuracy is 1.07% for unlearned labels and 84.77% for retained labels on average.
+For single-class tasks, it achieves complete unlearning with 0% accuracy for unlearned labels while maintaining 94.53% accuracy for retained labels.
+For multi-class tasks, the accuracy is 1.07% for unlearned labels and 84.77% for retained labels.
 Our framework accelerates unlearning by 99% compared to alternative methods.
-1. Introduction
-The explosive growth of data volume and the increasing complexity of applications necessitate models that can constantly update to adapt to new data, constraints, and standards (Jayaraman and Evans, 2019).
-Accordingly, the need for data privacy, the auditing of illegal information from models, and compliance with regulations and industry standards arises.
-This leads to the concept of “model forgetting” or “unlearning,” which involves removing specific information from trained models.
-The importance of unlearning lies in its ability to remove data that may raise privacy concerns, ensuring that the model can no longer recognize or rely on such data, in compliance with regulations such as the General Data Protection Regulation (GDPR) (Voigt and Von dem Bussche, 2017).
-Additionally, unlearning plays a crucial role in data correction, removing the impact of erroneous or biased data from the model and enhancing its accuracy and fairness.
-This not only protects user privacy but also boosts the model’s quality and reliability, aligning with both ethical standards and legal requirements.
-Machine unlearning becomes particularly challenging when models are consistently recorded with a directed acyclic graph (DAG) topology, where models inheriting from predecessors also require updates.
-Examples of such scenarios include a basic model being customized to meet various Machine Learning Operations (MLOps) lifecycle requirements (Matsui and Goya, 2023; Kreuzberger et al., 2023) or distilled to fit computational resource constraints (Polino et al., 2018).
-The concept of model inheritance can be applied to prevalent learning frameworks, such as Federated Learning (FL) (Li et al., 2020a), Distributed Data-Parallel Learning (DDPL) (Li et al., 2020c), Incremental Learning (IL) (Van de Ven et al., 2022), and Transfer Learning (TL) (Weiss et al., 2016).
-In this case, unlearning even a single model would need to unlearn the entire subgraph rooted at the origin to ensure all descendant models are consistent and relevant.
-Research Gap.
+I Introduction
+In the rapidly evolving landscape of data and application complexity, models require frequent updates to adapt to new data, constraints, and standards, highlighting the critical need for robust data privacy measures, auditing of illicit information, and adherence to regulatory and industry standards [jayaraman].
+The process of “model forgetting” or “unlearning” emerges as essential, involving the removal of specific information from trained models to address privacy concerns, ensuring models do not retain or utilize sensitive data in violation of regulations like the General Data Protection Regulation (GDPR) [voigt2017eu].
+Moreover, unlearning facilitates data correction by eliminating the influence of erroneous or biased inputs, thereby not only protecting user privacy but also enhancing the model’s accuracy, fairness, and overall quality.
+Many existing unlearning advances (e.g., [wang2025gru, yang2025exploring]) typically focus on removing the influence of specific data from a single model, without considering relationships among models.
+However, in an Inherited Model Network (IMN), where models are organized as a Directed Acyclic Graph (DAG), unlearning becomes significantly more complex: updates to one model must be consistently propagated to all its downstream inherited models.
+This structure introduces cascading effects and computational challenges that are unique to IMNs.
+This challenge arises in scenarios such as customizing a base model to meet various Machine Learning Operations (MLOps) lifecycle requirements [mbma] or distilling it to fit computational resource constraints [polino2018model].
+These situations can be abstracted into prevalent learning frameworks such as Federated Learning (FL) [li2020review], Distributed Data-Parallel Learning (DDPL) [ddap], Incremental Learning (IL) [IL], and Transfer Learning (TL) [TL].
+In such cases, unlearning even a single model necessitates unlearning the entire subgraph rooted at the origin to ensure that all descendant models remain consistent and relevant.
+Research gaps.
 Various machine unlearning techniques have been proposed.
-Unfortunately, the inheritance relationships among related models prevent most existing unlearning methods from being executed in parallel, including re-training (Kim and Woo, 2022; Liu et al., 2024b), gradient ascent (Zhang et al., 2024), or knowledge distillation methods (Sinha et al., 2023).
-On the other hand, Sharding, Isolation, Slicing, and Aggregated Training (SISA) (Bourtoule et al., 2021) is a commonly used efficient method that reduces the dependence between data by dividing it into multiple segments and training multiple models in parallel.
-However, SISA is also inapplicable when there are inheritance relationships between models (Liu et al., 2024a).
-This is due to the fact that SISA needs to be performed sequentially from the starting point of updating to the end of the subgraph, incurring significant overhead and resource consumption during unlearning.
-In this sense, there is a pressing need for an effective and efficient way to perform unlearning in the presence of model inheritance relationships.
-Research Question.
+Unfortunately, the inheritance relationships among related models prevent most existing unlearning methods from being executed in parallel, including re-training [retrain], gradient ascent [ga], or knowledge distillation methods [distil1Unl].
+On the other hand, Sharding, Isolation, Slicing, and Aggregated Training (SISA) [sisa, yu2024splitunlearning] is a commonly used efficient method that reduces the dependence between data by dividing it into multiple segments and training multiple models in parallel.
+However, SISA is also inapplicable when there are inheritance relationships between models [liu2024].
+This is because SISA needs to be performed sequentially from the starting point of updating to the end of the subgraph, incurring significant overhead and resource consumption during unlearning.
+In this sense, there is a pressing need for an efficient (i.e., unlearning speed) and effective (i.e., accuracy) way to perform unlearning in the presence of model inheritance relationships.
+Research questions.
 To address this research gap, we need to identify the commonalities across various prevalent frameworks in unlearning and determine how a new unlearning design can be universally applied.
-Specifically, we raise two key Research Questions (RQs):
-RQ1. How can unlearning requests be swiftly allocated to models that need updates and then prepare these models for unlearning tasks within a large-scale, interdependent model network?
+We raise two key Research Questions (RQs):
+RQ1. How can unlearning requests be quickly allocated to models that need updates and what unique features could be realized during unlearning in a large-scale and interdependent model network?
 RQ2. How can the model unlearning tasks be efficiently and effectively executed, especially when the unlearning of one model impacts subsequent models?
@@
 Contributions.
-This paper presents a novel parallel unlearning framework that can conduct efficient unlearning among large numbers of dependent models with complex inheritance relations, e.g., in FL, DDPL, IL, and TL.
-We first propose a new Unified Model Inheritance Graph (UMIG), to unveil the inheritance characteristic among models in various learning frameworks.
-With the assistance of the UMIG and Fisher Information Matrix (FIM), we develop the Fisher Inheritance Unlearning (FIUn) method that can efficiently pinpoint critical model parameters to be unlearned.
-This allows independent unlearning adjustments on large-scale inherited models.
-To further enable effective one-shot unlearning of knowledge inherited from multiple upstream, we design a merging FIM function that consolidates disparate FIMs and identifies the critical parameters that need updating across all sources.
-The contributions of this paper are summarized as follows:
-  • We propose the UMIG, which abstracts various prevalent learning frameworks, e.g., FL, DDPL, IL, and TL, in a unified fashion using DAGs to represent models’ interdependence.
-This facilitates the analysis of large-scale inheritance-oriented model networks.
-  • We design the FIUn method, which enables efficient parallel unlearning within the UMIG.
-The method leverages the FIM to quantify the significance of model parameters to unlearning tasks, and selectively scales the parameters according to their importance.
-  • We further develop the merging FIM function to consolidate the FIMs from multiple upstream unlearning models into a cohesive matrix, thereby aggregating the unlearning tasks inherited from those models.
-This design facilitates the one-shot removal of inherited knowledge and reduces computational overhead within the UMIG.
-Comprehensive experiments demonstrate that our method achieves complete unlearning for single-class unlearning tasks across all considered learning frameworks and model types, retaining an average accuracy of 94.53%.
-For multi-class unlearning tasks, the average unlearning accuracy drops to 1.07%, while the accuracy for retained labels reaches up to 84.77%.
-Additionally, our algorithm shows a 99% speed improvement across all considered frameworks compared to alternative methods.
-Paper Organization.
-The rest of the paper is organized as follows.
-Section 2 reviews the preliminaries.
-Section 3 provides the proposed UMIG, followed by introducing the novel FIUn method and merging FIM function in Section 4.
-Section 5 presents the experimental settings.
-Section 6 presents our comprehensive experimental results based on four learning frameworks.
-Section 7 reviews the related works.
-Section 8 concludes this work.
-2. Preliminaries
+This paper introduces a novel parallel unlearning framework for efficiently handling unlearning across numerous dependent models with complex inheritance relations, such as in FL, DDPL, IL, and TL.
+The contributions are summarized as follows:
+  • $\bullet$
+    We capture various unlearning scenarios in IMNs, which are chronologically DAG-based.
+By analyzing scenarios with single- and multi-root unlearning requests, and the impact of path depths and multi-path structures on unlearning, we provide precise and efficient identification of models to be unlearned.
+  • $\bullet$
+    We introduce the Fisher Inheritance Unlearning (FIUn) method to enable efficient parallel unlearning within DAGs.
+FIUn compares the Fisher Information Matrices (FIMs) of the initial unlearning models (upon the data to be unlearned) with the FIMs of all inherited models, identifies the model parameters that require updates, and then proportionally adjusts parameters in parallel across the inherited models based on their unlearning impact.
+  • $\bullet$
+    To handle multiple unlearning requests simultaneously, FIUn incorporates a Merging-FIM (MFIM) function that consolidates FIMs from multiple upstream models into a single unified matrix.
+This allows FIUn to address all unlearning scenarios, facilitating one-shot removal of inherited knowledge while greatly reducing computational overhead.
+Extensive experiments indicates that our method significantly outperforms all comparison methods in terms of unlearning speed on widely-accepted public image and text datasets.
+Compared to re-training, our method achieves comparable levels of unlearning accuracy.
+Even on large-scale models, our method demonstrates a certain degree of effectiveness and applicability.
+A preliminary version of this work has been posted on arXiv.
+[^1] This work is based on a preprint post: https://arxiv.org/abs/2408.08493.
+The rest of this paper is organized as follows.
+The related works are reviewed in Section II.
+The preliminaries are in Section III.
+The explored topological structure frameworks are discussed in Section IV.
+The designed parallel unlearning method in Section V.
+The experimental settings are presented in Section VI, followed by the evaluation in Section VII.
+Section VIII concludes this work.
+II Related Work
+Although FL [li2020federated], DDPL [ddap], IL [IL], and TL [TL] have addressed challenges such as data privacy [li2021survey], training efficiency [dean2012large], knowledge retention and unlearning [han2023anomaly], and cross-domain knowledge transfer [rajapaksha2023improving], systematic exploration combining these frameworks is limited.
+Most studies have only attempted simple method combinations without addressing their shared foundation model inheritance.
+This aspect is crucial in large-scale unlearning, where computational overhead and time consumption increase with inheritance depth, even when unlearning is parallelized at the same depth.
+This remains true across existing unlearning methods, which are discussed below.
+Federated Unlearning.
+Wang et al.
+[wang2024server] proposed a server-initiated federated unlearning framework, SIFU, which identifies and quantifies the impact of low-quality client data.
+By combining gradient ascent with high-quality data augmentation, the method efficiently eliminates the negative influence of such data on the global model.
+Tao et al.
+[tao2024communication] introduced the Total Variation (TV) Stability to measure model sensitivity to small data perturbations and modified the FedAvg algorithm to improve communication efficiency while enabling precise client- and sample-level unlearning.
+To address performance degradation caused by gradient conflicts, Pan et al.
+[pan2025federated] proposed the FedOSD method, which designs a dedicated unlearning loss function and optimized the model along the steepest descent direction closest to the target client’s gradient, without interfering with others, achieving a balance between unlearning and utility.
+Gao et al.
+[gao2024verifi] introduced the VeriFi framework, a framework that supports active verification of unlearning effects and includes a more efficient unlearning algorithm, uS2U, along with two non-invasive verification methods, vFM and vEM, enhancing the security and verifiability of federated unlearning.
+Distributed Data-Parallel Unlearning.
+In DDPL knowledge unlearning, handling the unlearning process is restricted to the data shard on the specific device, followed by model aggregation and iterative training.
+If unlearning is required from a specific shard, only the model trained on that shard is re-trained.
+Approximate machine unlearning directly modifies trained model parameters, employing methods such as 1) Gradient-based approaches [golatkar2020eternal], which update model gradients to gradually diminish the data’s influence, and 2) Data perturbation [tarun2023fast], which introduces noise to reduce specific data points’ impact on the model.
+These methods facilitate effective knowledge unlearning in DDPL.
+Incremental Unlearning.
+Research on unlearning in IL is still a developing area, but some recent methods show promise.
+Projected-Gradient Unlearning (PGU) [hoang2024learn] calculates gradients for the unlearning dataset, projects them onto the Core Gradient Space (CGS) derived from the training set, and subtracts the projected component.
+The residuals, now orthogonal to the CGS, are used to update model parameters.
+Effective for one-time and batch-wise unlearning, PGU is well-suited for incremental learning processes.
+Zuo et al.
+[zuo2024ecil] proposed a class-level unlearning method called eCIL-MU, which uses embedding to map data into vectors and store them in a vector database.
+This enables efficient and non-destructive unlearning by reusing the structure of class-incremental learning tasks.
+Transfer Unlearning.
+Research on unlearning specific target task knowledge has been relatively scarce in TL.
+To address this challenge, Sepahvand et al.
+[sepahvand2024data] proposed fine-tuning pre-trained models through data selection.
+Subsequently, the entire network was fine-tuned on the target task using gradient descent.
+Machine Unlearning.
+Nguyen et al.
+[nguyen2020variational] proposed a loss function minimizing Kullback-Leibler divergence to ensure certified data removal in Bayesian models, using two techniques to correct posterior inaccuracies and prevent catastrophic forgetting for small data subsets.
+Ginart et al.
+[ginart2019making] introduced Q-k-means and DC-k-means algorithms for k-means clustering, using quantized cluster centers and divide-and-conquer to efficiently update models without retraining, ensuring precise data removal while maintaining clustering quality.
+Kurmanji et al.
+[kurmanji2023towards] presented SCRUB, a teacher-student model optimizing student weights to diverge from the teacher’s output on the forget dataset ($D_{f}$) while aligning with the retain dataset ($D_{r}$), balancing forgetting and utility for bias removal, confusion resolution, and privacy.
+Zhu et al.
+[zhu2024decoupling] proposed TARF, using annealed gradient ascent and target-aware gradient descent to decouple specific concepts, preserving performance on unaffected data.
+Both modified model weights for efficient approximate unlearning without strict certified removal.
+Ours.
+Existing unlearning methods struggle with high computational costs, resource demands, and implementation challenges, particularly in large-scale settings.
+Our FIUn method enables fully parallel unlearning by adapting single-model techniques [foster2024fast, golatkar2020eternal] for IMNs, improving efficiency, resource utilization, and adaptability while achieving effectiveness comparable to re-training.
+III Preliminaries
 This section introduces the concepts, notation, and formulas of this paper, including machine unlearning and FIM.
-2.1. Machine Unlearning
-Machine unlearning refers to removing the influence of specific data subsets from a trained model while maintaining its performance on the remaining data (Foster et al., 2024a; Umer and Polikar, 2021).
-In other words, the goal is to eliminate the impact of specific labels on the model’s predictions while ensuring that the prediction accuracy for the remaining labels and the overall performance of the model remain unaffected.
-Let $D=\left\{\left(x_{i},y_{i}\right)\right\}^{N}_{i=1}$ represent a dataset where $x_{i}$ is the $i$-th training sample and $y_{i}\in\left\{1,\cdots,K\right\}$ is the corresponding class label.
-Let $C_{f}$ denote the set of labels to be unlearned, corresponding to the data set $D^{f}\subset D$ to be unlearned, and $C_{r}$ represent the set of retained labels, corresponding to the remaining data set $D^{r}=D\backslash D^{f}$.
-Unlearning aims to eliminate the impact of $D^{f}$ with labels in $C_{f}$ from the model while preserving its performance on $D^{r}$ with labels in $C_{r}$.
-In this context, we define $\phi_{\theta}\left(\cdot\right):X\rightarrow Y$ as a function parameterized by $\theta\in\mathbb{R}^{m}$, where $X\in\mathbb{R}^{n}$ and $Y\in\mathbb{R}^{K}$.
-The $k$-th component of $\phi_{\theta}\left(x\right)$ represents the probability that sample $x$ belongs to class $k$.
-The objective is to adjust $\theta$ such that the model effectively “forgets” the influence of $C_{f}$ while retaining accurate predictions on $C_{r}$.
-2.2. Fisher Information Matrix
-The FIM is a useful tool in statistics for assessing the accuracy of parameter estimates (Kirkpatrick et al., 2017; Golatkar et al., 2020).
+III-A Machine Unlearning
+Machine unlearning refers to removing the influence of specific data subsets from a trained model while maintaining its performance on the remaining data [foster2024zero].
+Let $D=\{(x_{i},y_{i})\}_{i=1}^{N}$ denote the full training dataset, where $x_{i}$ is the $i$-th input and $y_{i}\in\{1,\dots,K\}$ is its corresponding label.
+Let $D^{f}\subset D$ be the subset of samples to be unlearned (forget set), and $D^{r}=D\setminus D^{f}$ be the remaining samples (retain set).
+Let $\phi_{\theta}:X\rightarrow Y$ be the model parameterized by $\theta\in\mathbb{R}^{m}$, and $\phi_{\theta^{\prime}}$ be the updated model after unlearning.
+The unlearning process is considered successful if: (i) $\phi_{\theta^{\prime}}\perp D^{f}$, meaning that the predictions of $\phi_{\theta^{\prime}}$ are statistically independent of the removed data $D^{f}$, ideally resembling those of a model trained only on $D^{r}$; (ii) $L(\phi_{\theta^{\prime}},D^{r})\approx L(\phi_{\theta},D^{r})$, ensuring the model retains performance on the remaining data.
+In practice, this entails adjusting the model weights $\phi_{\theta}$ to obtain $\phi_{\theta^{\prime}}$ that removes the influence of $D^{f}$ while preserving knowledge from $D^{r}$.
+Crucially, model parameters are inherently heterogeneous—different weights contribute unequally to different classes, samples, or tasks [koh2017understanding].
+Effective unlearning therefore requires selective and targeted parameter adaptation.
+This motivates a range of gradient-based and sensitivity-aware approaches—including influence functions, Fisher Information, and other parameter importance metrics—that identify and update only the most relevant components in a principled and efficient manner.
+III-B Fisher Information Matrix
+FIM is a useful tool in statistics for assessing the accuracy of parameter estimates [golatkar2020eternal].
 It quantifies the impact of parameter changes on the probability distribution of observed data.
 Given a probability density function $p\left(x|\theta\right)$ conditioned on the model parameter $\theta$, an FIM $I\left(\theta\right)$ is the expected value of the second derivative of the negative log-likelihood function $\ell\left(\theta\right)$, as given by
-(1) $I\left(\theta\right)=\mathbb{E}\left[-\frac{\partial^{2}\ell\left(\theta\right)}{\partial\theta^{2}}\right],\ \text{where}\ \ell\left(\theta\right)=\ln p\left(x|\theta\right).$
-To simplify computation, in practice, the diagonal of the FIM can approximate the second derivative of the loss function (Steven, 1993; Aich, 2021).
-An equivalent form of $I\left(\theta\right)$ is given by
-(2) $I(\theta)=\mathbb{E}\left[\left(\left(\frac{\partial\ell\left(\theta\right)}{\partial\theta}\right)\left(\frac{\partial\ell\left(\theta\right)}{\partial\theta}\right)^{\top}\right)\right]\ .$
-In this paper, the FIM is employed to quantify the importance of data to the model parameters.
-By comparing the FIMs of the training dataset and the to-be-unlearned dataset, we can identify the critical parameters related to the knowledge that needs to be unlearned.
-3. Capturing Model-Inheritance: UMIG
-In this section, we explore the topological structure of four prevalent frameworks, FL, DDPL, IL, and TL, and analyze the impact of unlearning on the inherited models within these frameworks.
-We propose the UMIG, which can describe the model inheritance and update relationships in these learning frameworks in a unified manner.
-(a) FL vanilla and DDPL
-(b) Multi-layer/server FL
-(c) IL and TL
-(d) Decentralized FL
-Figure 1. These diagrams demonstrate various learning frameworks that can be constituted to the proposed UMIG.
-3.1. Diverse Learning Frameworks and Tasks
-3.1.1. Federated Learning
+$I\left(\theta\right)=\mathbb{E}\left[-\frac{\partial^{2}\ell\left(\theta\right)}{\partial\theta^{2}}\right],\ \text{where}\ \ell\left(\theta\right)=\ln p\left(x|\theta\right).$ (1)
+To simplify computation, in practice, the diagonal of FIM can approximate the second derivative of the loss function [steven1993fundamentals]. An equivalent form of $I\left(\theta\right)$ is given by
+$I(\theta)=\mathbb{E}\left[\left(\left(\frac{\partial\ell\left(\theta\right)}{\partial\theta}\right)\left(\frac{\partial\ell\left(\theta\right)}{\partial\theta}\right)^{\top}\right)\right]\ .$ (2)
+The FIM holds significant importance in the theory of machine unlearning.
+The natural gradient regards the FIM as a Riemannian metric tensor of the parameter space, used to define the optimal direction of parameter perturbation [amari1998natural].
+Furthermore, the FIM is the expected form of the Hessian, aligning with the concept of influence functions for analyzing the impact of individual samples [koh2017understanding].
+From an information-theoretic perspective, a larger FIM value indicates that the corresponding parameter carries more discriminative information about the training data and is therefore more critical to the model’s behavior.
+An intuitive baseline is to simply remove the columns in the last FC layer weights corresponding to the classes to be unlearned.
+However, this approach has significant limitations: (i) it is only applicable to class-level unlearning and cannot be extended to sample-level unlearning tasks; (ii) simple eliminating weight columns disrupts the decision boundaries of the retained data, leading to substantial accuracy degradation [lee2023undo]; and (iii) in IMNs, this method only affects the current node and fails to ensure consistent propagation of unlearning information across all inherited models.
+To address these issues, this paper employs the FIM to quantify the importance of model parameters to the training data.
+By comparing the FIMs of the training dataset and the unlearning dataset, we can identify the critical parameters related to the knowledge to be unlearned.
+IV Capturing Model-Inheritance
+In this section, we explore the topological structure of four prevalent frameworks, i.e., FL, DDPL, IL, and TL, and analyze the impact of unlearning on the inherited models within these frameworks.
+We propose using DAG, which can visualize and capture the model inheritance and update relationships in these learning frameworks in a unified manner.
+IV-A Diverse Learning Frameworks and Tasks
+Federated Learning.
 A central server creates a global model, which participants download to locally update the model parameters based on their local data.
@@
 The central server aggregates the parameters to update and redistribute the global model.
-This process repeats until the model converges (McMahan et al., 2017; Li et al., 2020b; Krauß et al., 2024).
-As shown in Figure 1(a), the topology of FL consists of nodes representing model states and edges representing parameter transmission and aggregation.
-Updates to model $A$ trigger updates in the global model $D$ and then to $E$, $F$, $G$, and $H$.
-FL also has other structures.
+This process repeats until the model converges [krauss2024automatic].
+As shown in Fig.
+1(a), the topology of FL consists of nodes representing model states and edges representing parameter transmission and aggregation.
+Updates to model $A$ trigger updates in the global model $D$ and then to $E$, $F$, $H$, and $I$.
+(a) FL vanilla and DDPL
+(b) Multi-layer/server FL
+(c) Decentralized FL
+(d) IL and TL
+Figure 1: These diagrams show learning frameworks modeled as a DAG, where one user may create multiple models, and multiple models may be created by different users.
 Multi-Layer FL.
-This variant introduces an intermediate aggregation layer to traditional FL (Chu et al., 2024; Li et al., 2021), as depicted in Figure 1(b).
+This variant introduces an intermediate aggregation layer to traditional FL [li2021survey], as depicted in Fig.
+1(b).
 This approach adds multiple servers that perform intermediate aggregation between the clients and the global server, thus reducing the burden on the global server, improving system scalability, and enhancing model training.
-DAG-FL.
-As shown in Figure 1(d), multiple clients update and aggregate models hierarchically in DAG-FL (Cao et al., 2021a, b; Yu et al., 2023).
+Decenralized FL.
+As shown in Fig.
+1(c), multiple clients update and aggregate models hierarchically in DAG-FL [yu2023ironforge].
 Each node can receive model updates from multiple parent nodes and send the updated results to multiple child nodes.
 This structure enables efficient model training and more flexible communication, better adapting to complex network environments and different application needs.
-3.1.2. Distributed Data-Parallel Learning
-Training tasks are typically segmented into multiple sub-tasks, distributed among various computing nodes or servers for execution, as documented in (Ge et al., 2023; Dean et al., 2012; Abadi et al., 2016; Paszke et al., 2019; Li et al., 2014).
-Each node independently processes its designated sub-task and periodically synchronizes its model parameters with others to collectively refine the global model.
-This iterative process is crucial, particularly in DDPL topologies — similar to the FL topology depicted in Figure 1(a) — which require the sub-tasks $E$, $F$, and $H$, as well as the aggregated model $I$, to be updated continually with each new iteration of the global model $D$.
-3.1.3. Incremental Learning
-Incremental (or “Continual”) learning entails training a model initially on a dataset and then progressively updating it as new data arrives, ensuring adaptability to the latest information (Han et al., 2023; Rajapaksha et al., 2023).
-IL can be visualized as shown in Figure 1(c), where nodes represent different updated states of the model, and edges depict the transition between model states after data updates.
+Distributed Data-Parallel Learning.
+Training tasks are divided into sub-tasks and distributed across computing nodes or servers [liu2024fishers, dean2012large].
+Each node processes its sub-task independently and periodically synchronizes parameters to update the global model.
+This iterative process is essential in DDPL topologies—similar to the FL topology in Fig.
+1(a)—where sub-tasks $E$, $F$, and $H$ and the aggregated model $I$ must be updated with every new iteration of the global model $D$.
+Incremental Learning.
+IL entails training a model initially on a dataset and then progressively updating it as new data arrives, ensuring adaptability to the latest information [han2023anomaly].
+IL can be visualized as shown in Fig.
+1(d), where nodes represent different updated states of the model, and edges depict the transition between model states after data updates.
 If the initial state model $G$ continues to update, then the subsequent models $A$ and $B$ that receive new data, as well as their related models, also need updates.
-3.1.4. Transfer Learning
-TL involves pre-training a base model on one dataset and then applying this model to a new task.
-Typically, this process includes fine-tuning the parameters of the last few layers to better suit the new task (Sendner et al., 2023; Rajapaksha et al., 2023).
-The model continues to train on the specific dataset of the new task to adapt fully to its requirements.
-Through ongoing fine-tuning and training, the model is optimized for the best performance on the new task.
-The TL topology, similar to the IL topology shown in Figure 1(c), also follows a comparable model updating process.
-In these diverse learning frameworks, the inheritance relationship between models is pivotal.
-When a base model is updated or optimized, all corresponding child models derived from it must also be updated.
-To accurately capture this inheritance relationship, we employ the DAG topology as an effective abstraction method to capture the dependencies among models in these learning frameworks.
-The use of the DAG ensures that when a single model is updated, the entire subgraph—originating from this model and encompassing all related child models—is appropriately updated.
-3.2. Unified Model Inheritance Graph (UMIG)
-Figure 2. The darker blue nodes are all model nodes inheriting from the starting nodes $n_{s}$ and $n_{f}$ in lighter blue. A model node $n_{\ast}$ corresponds to a model $w_{\ast}$.
-To uniformly handle and analyze the inheritance and update relationships of models across various learning frameworks, including FL, DDPL, IL, and TL, we propose a DAG-based UMIG.
-It provides a comprehensive framework to represent components such as models, data, and tasks and their interdependencies within these frameworks.
-It also facilitates a structured and intuitive understanding of these complex model networks.
-Figure 2 illustrates the ability of UMIG to depict the inheritance relationships between models.
-As illustrated in Figure 2, we define a weighted directional graph $\mathcal{G}=(N,E)$.
-Herein, $N=\left\{n_{1},n_{2},\ldots,n_{m}\right\}$ represents the model nodes in the network, more specifically, a set of model parameters in a specific state or task.
-Moreover, $E=\left\{e_{01},e_{02},\ldots,e_{mm}\right\}$ represents the inheritance relationships between model nodes, encompassing parameter transfer and task inheritance.
-For instance, in FL and DDPL, the edges may represent the transfer of parameters from local to global models.
-By contrast, in TL, the edges signify the inheritance of tasks from a pre-trained model to a fine-tuned model.
+Transfer Learning.
+TL pre-trains a base model on one dataset and adapts it to a new task by fine-tuning, typically on the last layers, using task-specific data [rajapaksha2023improving].
+Through continued fine-tuning, the model aligns with the new task and achieves optimal performance.
+The TL topology mirrors the IL topology in Fig.
+1(d), following a similar update process.
+In these learning frameworks, inheritance between models is crucial: when a base model is updated, all its child models must also be updated.
+We use a DAG topology to represent these dependencies, ensuring that updates to one model propagate through its entire subgraph of descendants.
+IV-B Inherited Model Networks
+Figure 2: The dark blue nodes are all model nodes inheriting from the starting nodes $n_{s}$ and $n_{f}$ in light blue.
+A model node $n_{\ast}$ corresponds to a model $w_{\ast}$.
+$w_{s}$ and $w_{f}$ can originate from different users, with the possibility of simultaneous unlearning requests.
+We use a DAG to represent and analyze model inheritance and updates across frameworks like FL, DDPL, IL, and TL.
+Time is integrated into the representation: each node corresponds to a publication moment, and edges follow the chronological order from left to right.
+This captures both inheritance dependencies and the temporal sequence of model publication, providing a clear view of complex model networks.
+As shown in Fig.
+2, we model the system as a weighted directed graph $\mathcal{G}=(N,E)$.
+The nodes $N={n_{1},n_{2},\ldots,n_{m}}$ represent model states or tasks, each corresponding to a set of parameters.
+The edges $E={e_{ij}}$ capture inheritance relations, such as parameter transfer or task inheritance.
+In FL and DDPL, edges may denote parameter transfer from local to global models, while in TL they represent task inheritance from pre-trained to fine-tuned models.
+Edge direction is determined by temporal order, reflecting when nodes are published, and by reference relations, indicating when one node explicitly refers to another.
 Based on the source and state of the model, we further categorize the nodes into two types:
-  • Discovery Nodes refer to models trained on data with an unlearned label set, enabling these nodes to adapt to the updated classification ability within the corresponding subgraph of the DAG.
-The root model nodes $n_{s}$ and $n_{f}$ in Figure 2 belong to this type.
-  • Inherited Nodes comprise the remaining nodes in the subgraph, except the discovery nodes.
-Inherited nodes point to the discovery or other inherited nodes, and focus on refining classification performance.
-The model nodes $n_{h},n_{j},n_{l},n_{p},n_{g},n_{i},n_{n}$ belong to this type.
-3.3. UMIG-Assisted Unlearning
-With an emphasis on knowledge inheritance and unlearning tasks, UMIG allows a model to retain and utilize previously learned useful information when receiving new data, achieving an effective combination of old and new knowledge.
-UMIG also allows the model to remove specific data for continuous optimization and performance improvement.
-The UMIG-assisted unlearning consists primarily of the following two processes:
+  • $\bullet$
+    Discovery Nodes are models trained on data with unlearned labels, allowing adaptation to updated classification capabilities within their DAG subgraph.
+Root model nodes $n_{s}$ and $n_{f}$ exemplify this type in Fig.
+2.
+  • $\bullet$
+    Inherited Nodes are all non-discovery nodes in the subgraph that connect to discovery or other inherited nodes, focusing on classification refinement.
+Nodes $n_{h},n_{j},n_{l},n_{p},n_{g},n_{i},n_{n}$ belong to this type.
+IV-C DAG-Assisted Unlearning
+With an emphasis on knowledge inheritance and unlearning tasks, DAG allows a model to retain and utilize previously learned useful information when receiving new data, achieving an effective combination of old and new knowledge.
+DAG also allows the model to remove specific data for continuous optimization and performance improvement.
+The DAG-assisted unlearning consists primarily of the following two processes:
+  • $\bullet$
     Locating unlearning subgraphs.
-Based on the data that needs to be unlearned, a breadth-first search is conducted to identify the discovery nodes, which then spawn the subgraphs associated with the discovery nodes.
+When a user wants to unlearn certain data, an unlearn request is initiated.
+The system first identifies the nodes related to that user, which are referred to as discovery nodes, which then spawn the subgraphs associated with the discovery nodes.
+  • $\bullet$
     Updating models within unlearning subgraphs.
 All model nodes, including discovery and inherited nodes within the identified unlearning subgraphs, are updated to unlearn the parts of their models affected by the data.
-By performing a breadth-first search of the entire model network, one or more discovery nodes are identified based on unlearning requests related to specific data.
 These discovery nodes serve as the roots for the unlearning subgraphs.
-As depicted in Figure 2, the unlearning subgraphs may overlap since inherited nodes may inherit from multiple discovery nodes.
+As depicted in Fig.
+2, the unlearning subgraphs may overlap since inherited nodes may inherit from multiple discovery nodes.
 For example, the discovery nodes, $n_{s}$ and $n_{f}$, lead to two overlapping subgraphs with inherited nodes.
@@
 Each path represents a route of knowledge transfer via iteration, and the number of paths and differences in depth determine the complexity of the inheritance relationship.
-A complex structure emerges when multiple paths are at varying depths from an inherited node to its corresponding discovery node(s), e.g., $n_{f},n_{g},n_{i}$ or $n_{s},n_{f},n_{h},n_{j}$.
+A complex structure emerges when multiple paths are at varying depths from an inherited node to its corresponding discovery node(s), e.g., $n_{f},n_{g},n_{i}$ or $n_{s},n_{h},n_{j}$.
 If unlearning operations are conducted simultaneously in terms of depth (in other words, nodes at the same depth from the root(s) are processed simultaneously and rely on sequential dependencies), nodes such as $n_{i}$ and $n_{j}$ may require multiple updates.
@@
 In cases where unlearning is conducted asynchronously, regardless of depth, both scenarios may require multiple updates.
-Answer to RQ1: We propose the UMIG to effectively visualize model interdependence and capture the model inheritance in a structured way, leveraging the DAG structure.
-Assisted by UMIG, the models inheriting the knowledge to be removed can be rapidly identified.
-Moreover, the UMIG offers a systematic way to collect and prepare inheritance information, thus facilitating the unlearning process of inherited models.
-We also examine various learning frameworks of inherited models based on the numbers of their upstream models and the paths of knowledge inheritance to elucidate the emerging challenges and considerations of the unlearning process in inherited nodes due to model inheritance.
-Figure 3.
-The unlearning process in $\mathcal{G}$ consists of five steps: 1) Preparation.
-The discovery nodes $\{n^{\scriptscriptstyle D}_{i}\}$ are identified by performing a breadth-first search, with unlearning data $\hat{D}^{\scriptscriptstyle D}_{i}$ containing one or more labels.
+Answer to RQ1: We use a DAG to visualize model inheritance.
+Each node marks a specific time, and edges follow the chronological order of model publication, giving a clear view of complex inheritance networks.
+The DAG also helps quickly identify models whose knowledge should be removed and systematically collects inheritance information to simplify unlearning.
+We analyze different learning frameworks, focusing on upstream models and inheritance paths to capture key unlearning scenarios: single-root, multi-root, balanced-path, and imbalanced-path.
+Figure 3: The unlearning process in $\mathcal{G}$ consists of five steps: 1) Preparation.
+The discovery nodes $\{n^{\scriptscriptstyle D}_{i}\}$ are identified, with unlearning data $\hat{D}^{\scriptscriptstyle D}_{i}$ containing one or more labels.
 Subsequently, the unlearning graph $\mathcal{\hat{G}}$ is located which root at $\{n^{\scriptscriptstyle D}_{i}\}$.
@@
 5) Update $n_{j}$ by comparing the merged unlearning FIM $\hat{F}_{j}^{\scriptscriptstyle M}$ and $F_{j}$.
-Steps 2, 3, 4 and 5 apply to all nodes in $\mathcal{\hat{G}}$
-4. FIUn-based Parallel Unlearning
-This section introduces a novel parallel unlearning method, Fisher Inheritance Unlearning (FIUn), designed to address the challenges associated with machine unlearning in large-scale UMIG model networks.
+Steps 2, 3, 4 and 5 apply to all nodes in $\mathcal{\hat{G}}$ .
+V FIUn-Based Parallel Unlearning
+This section introduces a novel parallel unlearning method, Fisher Inheritance Unlearning (FIUn), designed to address the challenges associated with machine unlearning in large-scale DAG model networks.
 A trusted learning network is considered where all learning contributors follow the designed unlearning process.
-The FIUn method operates under the general expression provided by the UMIG and leverages FIMs.
+The FIUn method operates under the general expression provided by the DAG and leverages FIMs.
 It initiates the unlearning process by identifying specific model parameters that need updates, making the process more systematic.
@@
 Notably, the FIUn method allows for the independent application of unlearning steps to each impacted model, facilitating a swift and effective parallel unlearning across extensive model networks.
-The FIUn method is depicted in Figure 3, and the procedural details are given in Algorithm 1.
+The FIUn method is depicted in Fig.
+3, and the procedural details are given in Algorithm 1.
 1
@@
 3Parameter: $\gamma$, $\tau$, $\eta$
-4Use breadth-first search to identify discovery nodes $\{n^{\scriptscriptstyle D}_{i}\}$ and consequential unlearning graph $\mathcal{\hat{G}}$
+4Identify discovery nodes $n^{\scriptscriptstyle D}_{i}$} and the resulting unlearning graph $\mathcal{\hat{G}}$
 5 for $n^{\scriptscriptstyle D}_{i}$ in $\{n^{\scriptscriptstyle D}_{i}\}$ parallel do
-6       Calculate unlearning FIM $\hat{F}^{\scriptscriptstyle D}_{i}$ with Eq. (3)
+6  Calculate unlearning FIM $\hat{F}^{\scriptscriptstyle D}_{i}$ with (3)
 7 end for
@@
 9for $n_{j}$ in $\mathcal{\hat{G}}$ parallel do
-10       Calculate model FIM $F_{j}$ with Eq. (4)
-11       Merge unlearning FIMs $\{\hat{F}^{\scriptscriptstyle D}_{j,k}\}$ with Eq. (5)
-12       Update model parameters of node $n_{j}$ with Eq. (6)
+10  Calculate model FIM $F_{j}$ with (4)
+11  Merge unlearning FIMs $\{\hat{F}^{\scriptscriptstyle D}_{j,k}\}$ with (5)
+12  Update model parameters of node $n_{j}$ with (6)
 13 end for
@@
 Algorithm 1 Fisher Inheritance Unlearning (FIUn)
-4.1. Preparation
-Upon receiving an unlearning task $\mathcal{T}$, accompanied by the unlearning dataset $\hat{D}$, FIUn initiates by conducting a breadth-first search from the starting node on the model network $\mathcal{G}$ to identify the discovery nodes $\{n^{\scriptscriptstyle D}_{i}\}$.
-The discovery nodes are the first nodes exhibiting significant recognition of the unlearning data, with $n^{\scriptscriptstyle D}_{i}$ representing the $i$-th discovery node.
-Subsequently, FIUn delineates the unlearning graph $\mathcal{\hat{G}}$, which is rooted at the identified discovery nodes.
-FIUn then updates all nodes in the unlearning graph to unlearn the requested unlearning knowledge.
-4.2. FIUn
+V-A FIUn
 The main idea of FIUn is to identify critical model parameters by comparing the FIM from the unlearning dataset with the FIM from the training dataset.
@@
 The FIUn process unfolds as follows.
-4.2.1. Calculate unlearning FIMs of discovery nodes
-The FIUn method initiates by assessing the impact of the unlearning data on the model parameters of the discovery nodes, utilizing the FIM as specified in Eq.
-(3).
+Preparation.
+Upon receiving an unlearning task $\mathcal{T}$ initiated by a user, FIUn first identifies the set of discovery nodes $\{n^{\scriptscriptstyle D}_{i}\}$ related to that user from the model network $\mathcal{G}$, with $n^{\scriptscriptstyle D}_{i}$ representing the $i$-th discovery node.
+Subsequently, FIUn delineates the unlearning graph $\mathcal{\hat{G}}$, which is rooted at the identified discovery nodes.
+FIUn then updates all nodes in the unlearning graph to unlearn the requested unlearning knowledge.
+Step-1: Calculate unlearning FIMs of discovery nodes.
+The FIUn method initiates by assessing the impact of the unlearning data on the model parameters of the discovery nodes, utilizing the FIM as specified in (3).
 This assessment is conducted using the model parameters and the unlearning data (see line 5, Algorithm 1).
 The unlearning FIM for the $i$-th identified discovery node, $n^{\scriptscriptstyle D}_{i}$, is denoted by $\hat{F}^{\scriptscriptstyle D}_{i}$ and is calculated as
-(3) $\hat{F}^{\scriptscriptstyle D}_{i}\!\!=\!\mathbb{E}\!\left[\!\left(\!\!\left(\frac{\partial\ln p(\hat{D}_{i}^{\scriptscriptstyle D}|w^{\scriptscriptstyle L}_{i})}{\partial w^{\scriptscriptstyle L}_{i}}\right)\!\!\!\left(\frac{\partial\ln p(\hat{D}_{i}^{\scriptscriptstyle D}|w^{\scriptscriptstyle L}_{i})}{\partial w^{\scriptscriptstyle L}_{i}}\!\right)^{\!\!\!\intercal}\right)\!\!\Bigg{|}_{w^{\ast}_{\hat{D}_{i}^{\scriptscriptstyle D}}}\!\right],$
+$\hat{F}^{\scriptscriptstyle D}_{i}\!\!=\!\mathbb{E}\!\left[\!\left(\left(\frac{\partial\ln p(\hat{D}_{i}^{\scriptscriptstyle D}|w^{\scriptscriptstyle L}_{i})}{\partial w^{\scriptscriptstyle L}_{i}}\right)\left(\frac{\partial\ln p(\hat{D}_{i}^{\scriptscriptstyle D}|w^{\scriptscriptstyle L}_{i})}{\partial w^{\scriptscriptstyle L}_{i}}\!\right)^{\intercal}\right)\Bigg|_{w^{\ast}_{\hat{D}_{i}^{\scriptscriptstyle D}}}\!\right],$ (3)
 where $\hat{D}^{\scriptscriptstyle D}_{i}$ represents the unlearning data for $n^{\scriptscriptstyle D}_{i}$, $w^{\scriptscriptstyle L}_{i}$ denotes the last layer parameters of the model $w_{i}$ at node $n^{\scriptscriptstyle D}_{i}$, and $w^{\ast}_{\hat{D}_{i}^{\scriptscriptstyle D}}$ refers to the optimal parameters learned on $\hat{D}_{i}^{\scriptscriptstyle D}$.
-The unlearning FIM highlights the importance of each parameter in the last layer with respect to the specific unlearning task (Lee et al., 2023; Zhang et al., 2021).
-In FIUn, the FIM calculation is restricted to the last layer parameters while other layers remain frozen, aimed at reducing the computational complexity associated with FIM calculations, which is a challenge in traditional FIM-based unlearning methods.
-Using the last layer for unlearning can focus on the most relevant features.
-Since the last layer directly affects the final decision of the model, adjusting its parameters can effectively facilitate unlearning goals while maintaining stable performance on data not subject to unlearning (Kim et al., 2024).
-4.2.2. Calculate model FIMs of all nodes in unlearning graph
+The unlearning FIM highlights the importance of last-layer parameters for specific unlearning tasks [lee2023layer].
+In FIUn, FIM calculations are restricted to the last layer while freezing other layers, thereby reducing the computational complexity inherent in traditional FIM-based methods.
+Focusing on the last layer targets the most relevant features, as it directly influences the model’s final decision.
+This approach effectively facilitates unlearning goals while maintaining stable performance on unaffected data.
+Step-2: Calculate model FIMs of all nodes in the unlearning graph.
 The FIUn method proceeds by calculating the model FIMs of all models within the unlearning graph to identify the impact of model parameters in each node (see line 8, Algorithm 1).
-For the $j$-th node $n_{j}$ in the unlearning graph $\mathcal{\hat{G}}$, either a discovery node or an inherited node, the model FIM, denoted by $F_{j}$, is given by
-(4) $F_{j}\!=\!\mathbb{E}\!\left[\!\left(\!\!\left(\frac{\partial\ln p(D_{j}|w^{\scriptscriptstyle L}_{j})}{\partial w^{\scriptscriptstyle L}_{j}}\right)\!\!\!\left(\frac{\partial\ln p(D_{j}|w^{\scriptscriptstyle L}_{j})}{\partial w^{\scriptscriptstyle L}_{j}}\right)^{\!\!\!\intercal}\right)\!\Bigg{|}_{w^{\ast}_{D_{j}}}\right]\!,$
+For any node $n_{j}$ in the unlearning graph $\mathcal{\hat{G}}$, the model FIM $F_{j}$ is given by
+$F_{j}\!=\!\mathbb{E}\!\left[\!\left(\left(\frac{\partial\ln p(D_{j}|w^{\scriptscriptstyle L}_{j})}{\partial w^{\scriptscriptstyle L}_{j}}\right)\left(\frac{\partial\ln p(D_{j}|w^{\scriptscriptstyle L}_{j})}{\partial w^{\scriptscriptstyle L}_{j}}\right)^{\intercal}\right)\!\Bigg|_{w^{\ast}_{D_{j}}}\right]\!,$ (4)
 where $D_{j}$ is the training data for $n_{j}$, $w^{\scriptscriptstyle L}_{j}$ denotes the last layer parameters of the model $w_{j}$ at node $n_{j}$, and $w^{\ast}_{D_{j}}$ denotes the optimal parameters learned on $D_{j}$.
@@
 Similar to the unlearning FIM, only the FIM for the last layer parameter is calculated, aiming to minimize computational complexity.
-4.2.3. Merge unlearning FIMs for all nodes in the unlearning graph
+Step-3: Merge unlearning FIMs for all nodes in the unlearning graph.
 The FIUn method estimates the unlearning FIM for all nodes within the unlearning graph, including both discovery and inherited nodes (see line 9, Algorithm 1).
 The unlearning FIMs from discovery nodes are merged to form a collective unlearning FIM for each node in the unlearning graph.
-This merger is crucial as the linkages between the unlearning data and parameters are embedded within the model and can persist through model inheritance (Foster et al., 2024b), especially in multi-root scenarios where an inherited node may trace back to several discovery nodes.
-A novel merging FIM function is introduced, denoted by $\Phi\left(\cdot\right)$, and is defined as
-(5) $\hat{F}^{\scriptscriptstyle M}_{j}=\Phi\left(\{\hat{F}^{\scriptscriptstyle D}_{j,k}\}\right),$
+This merger is crucial as the linkages between the unlearning data and parameters are embedded within the model and can persist through model inheritance [foster2024fast], especially in multi-root scenarios of DAG, where an inherited node may trace back to several discovery nodes.
+Sequential FIM calculation of the unlearned data from multiple roots is required for existing FIM methods [foster2024fast, golatkar2020eternal], but this significantly reduces efficiency in IMNs.
+To address this, a novel Merging-FIM (MFIM) function is introduced specifically for this context, denoted by $\Phi\left(\cdot\right)$, and is defined as
+$\hat{F}^{\scriptscriptstyle M}_{j}=\Phi\left(\{\hat{F}^{\scriptscriptstyle D}_{j,k}\}\right),$ (5)
 where $\hat{F}^{\scriptscriptstyle M}_{j}$ is the merged unlearning FIM for the $j$-th node in the unlearning graph, the set $\{\hat{F}^{\scriptscriptstyle D}_{j,k}\}$ collects the unlearning FIMs of discovery nodes within $\mathcal{\hat{G}}$ that can be traced back from node $n_{j}$, and $k$ denotes the $k$-th discovery node reachable from $n_{j}$.
-In this paper, the element-wise maximum is employed for the merging FIM function.
+In this paper, the element-wise maximum is employed for the MFIM function.
 Specifically, $\hat{F}^{\scriptscriptstyle M}_{j}$ assembles the largest elements among the unlearning FIMs of discovery nodes in multi-root scenarios and takes the only unlearning FIM in single-root scenarios.
-Figure 4.
-There are three discovery nodes that need to undergo the unlearning process, namely model $w_{j}$, model $w_{s}$ and model $w_{f}$.
+Figure 4: Three discovery nodes need to undergo the unlearning process, namely model $w_{j}$, model $w_{s}$ and model $w_{f}$.
 The labels to be unlearned are label 1 and label 2.
@@
 Model $w_{h}$ inherits from these three discovery nodes, so we merge the unlearning FIMs of the discovery nodes and perform the unlearning of label 1 and label 2 on model $w_{h}$ together.
-As shown in Figure 4, the model $w_{h}$ inherits from models $w_{j}$, $w_{s}$, and $w_{f}$ and needs to remove all requested unlearning data, though each discovery node contains only a portion of the unlearning data.
+As shown in Fig.
+4, the model $w_{h}$ inherits from models $w_{j}$, $w_{s}$, and $w_{f}$ and needs to remove requested unlearning data, though each discovery node contains only a portion of the unlearning data.
 The FIM of each model reflects the sensitivity of its parameters to the unlearning data, which can be interpreted as parameter importance.
-By taking the maximum of the corresponding elements across these unlearning FIMs, the merged FIM encompasses all unlearning tasks from the discovery nodes and effectively identifies the parameters that need updating to achieve complete unlearning.
-4.2.4. Update models in unlearning graph
+By taking the maximum of the corresponding elements across these unlearning FIMs, the merged FIM encompasses all unlearning tasks from the discovery nodes and identifies the parameters that need updating to achieve complete unlearning in a single process.
+Step-4: Update models in the unlearning graph.
 In the final step, the FIUn method identifies which parameters need to be updated by comparing the merged FIM with the model FIM, as described in line 10 of Algorithm 1.
-The outcomes of this comparison are also utilized to scale the parameters.
+The comparison, based on the Selective Synaptic Dampening (SSD) [foster2024fast], is also utilized to scale the parameters.
 For node $n_{j}$ in the unlearning graph, the parameters are updated as
-(6) $\hat{w}^{\scriptscriptstyle L}_{j,l}=\begin{cases}\min(\tau\frac{F_{j,l}}{\hat{F}^{\scriptscriptstyle M}_{j,l}},\eta)w^{\scriptscriptstyle L}_{j,l},&\text{if }\frac{\hat{F}^{\scriptscriptstyle M}_{j,l}}{F_{j,l}}>\gamma;\\ w^{\scriptscriptstyle L}_{j,l},&\text{if}\ \frac{\hat{F}^{\scriptscriptstyle M}_{j,l}}{F_{j,l}}\leq\gamma,\end{cases}$
+$\hat{w}^{\scriptscriptstyle L}_{j,l}=\begin{cases}\min(\tau\frac{F_{j,l}}{\hat{F}^{\scriptscriptstyle M}_{j,l}},\eta)w^{\scriptscriptstyle L}_{j,l},&\text{if }\frac{\hat{F}^{\scriptscriptstyle M}_{j,l}}{F_{j,l}}>\gamma;\\ w^{\scriptscriptstyle L}_{j,l},&\text{if}\ \frac{\hat{F}^{\scriptscriptstyle M}_{j,l}}{F_{j,l}}\leq\gamma,\end{cases}$ (6)
 where $w^{\scriptscriptstyle L}_{j,l}$ and $\hat{w}^{\scriptscriptstyle L}_{j,l}$ are the $l$-th parameter in the last layer before and after the unlearning update, respectively.
@@
 Non-critical parameters are left unchanged to maintain accuracy on data that does not require unlearning.
-4.3. Discussion
-From the perspective of the FIM, the impact of data on specific model parameters is consistently localized at precise positions within the parameter space (Foster et al., 2024b).
+Specific datasets can also be retained by extending the formulations in (5) and (6).
+In this setting, the unlearning FIM is replaced with a retention FIM, and the role of the parameter $\gamma$ shifts from suppressing to amplifying the weights of critical parameters.
+Since larger FIM values indicate that the corresponding parameters carry more discriminative information about the training data, we adjust these key parameters proportionally, thereby strengthening the model’s memory and dependence on the retained dataset.
+V-B Discussion
+From the perspective of the FIM, the impact of data on specific model parameters is consistently localized at precise positions within the parameter space [foster2024fast].
 The stable localization of the FIM across iterations helps precisely locate the parameter locations for unlearning at the class, client, and sample levels using data tied to specific label sets.
-By fine-tuning $\tau$, $\eta$, and $\gamma$ in Eq.
-(6) to match the sample size of the class being unlearned, the task transits flexibly between the class-level (Foster et al., 2024b) and sample-level unlearning (Golatkar et al., 2020).
+By fine-tuning $\tau$, $\eta$, and $\gamma$ in (6) to match the sample size of the class being unlearned, the task transits flexibly between the class-level [foster2024fast] and sample-level unlearning [golatkar2020eternal].
 Specifically, when the unlearning FIM mirrors the model’s FIM with a class-level setting, it becomes a client-level unlearning task.
@@
 Consequently, employing FIUn disrupts the sequential dependencies traditionally required for unlearning, thus enabling fully parallel unlearning processes across any number of subgraphs.
-Not only can models across different subgraphs undergo parallel unlearning, but those within the same subgraph at various depths can also be unlearned in parallel, thanks to a feature referred to as Hyper-Distance in this paper.
-This Hyper-Distance is achieved without compromising the integrity of the inheritance relationships.
-Hyper-Distance ensures that both unlearned and retained knowledge is effectively managed along the inheritance paths.
-This efficiency is possible because the FIUn method utilizes FIMs for unlearning, enabling a class-wise inheritance approach.
-Each model effectively incorporates a union of label sets from its predecessors, ensuring that the inheritance is preserved.
-Simultaneously, by focusing on label differences, we achieve fully parallel unlearning—making depths and paths irrelevant to the process.
-Answer to RQ2: The FIUn method enables independently removing knowledge from each inherited model.
+Models across subgraphs and at different depths within the same subgraph can undergo parallel unlearning, enabled by Hyper-Distance, where learned knowledge’s parameter impact is independent of model inheritance depth.
+Each model integrates predecessors’ label sets to preserve inheritance; focusing on label differences achieves full parallel unlearning, making depths/paths irrelevant without compromising inheritance integrity.
+This is possible because the FIUn method utilizes FIMs for unlearning, enabling a class-wise inheritance approach.
+Answer to RQ2: The FIUn method enables the independent removal of knowledge from each inherited model.
 This method utilizes the FIM to obtain the Hyper-Distance property which breaks sequential dependencies among models, enabling fully parallel unlearning on inherited models and significantly decreasing the execution time of unlearning tasks.
-To address the complexity arising from multiple upstream models, the merging FIM function is developed to aggregate unlearning FIMs from those models, facilitating the efficient, one-shot removal of inherited knowledge.
-We also map the FIUn method and merging FIM function to the UMIG-assisted unlearning scenarios described in Section 3.3.
-This demonstrates how the proposed approach efficiently facilitates unlearning through the property of Hyper-Distance.
+To address the complexity arising from multiple upstream models, the MFIM function is developed to aggregate unlearning FIMs from those models, facilitating the efficient, one-shot removal of inherited knowledge.
+We also map the FIUn method and the MFIM function to the DAG-assisted unlearning scenarios described in Sec.
+IV-C.
+This shows how the proposed method efficiently facilitates unlearning through the property of Hyper-Distance.
 Converting Imbalance into Balance.
-This scenario is particularly significant due to the fact that existing unlearning methods suffer from inevitable sequential dependencies.
-FIUn leverages the Hyper-Distance property to circumvent the influences of depths and paths within the inheritance topology.
-Consequently, imbalanced-path scenarios can be managed as effectively as balanced-path scenarios, where only one update operation is needed, regardless of whether unlearning is conducted synchronously or asynchronously, i.e., only one unlearning FIM needs to be computed for each inherited node.
+This scenario is particularly significant since existing unlearning methods face unavoidable sequential dependencies.
+FIUn utilizes the Hyper-Distance property to bypass the effects of depths and paths in the inheritance topology.
+As a result, imbalanced-path scenarios are handled as efficiently as balanced ones, requiring just one update operation per inherited node, regardless of whether unlearning is synchronous or asynchronous.
 Converting Multi-root into Single-root.
-The merging FIM function is proposed to refine multi-root scenarios.
-This function merges the FIMs of the corresponding unlearned label sets from multiple discovery nodes into a single FIM, as shown in Figure 4.
+The MFIM function is proposed to refine multi-root scenarios.
+This function merges the FIMs of the corresponding unlearned label sets from multiple discovery nodes into a single FIM, as shown in Fig.
+4.
 Specifically, overlapping inherited nodes require only one update operation to eliminate the impact of multiple discovery nodes by conducting a parameter-wise merger, where only the most impactful parameters take effect in this FIM.
 This function eliminates the need for multiple comparisons among various unlearning FIMs in multi-root scenarios, hence enhancing unlearning efficiency.
-4.4. Implementation
-4.4.1. System Setting
-In practice, FIUn can be handled differently depending on the system architecture, as described in the following:
-  • Centralized: Model FIMs are computed and attached along with the model being published and sent to the central coordinator.
-When an unlearning request is initiated, the central server performs unlearning by merging the unlearning FIMs of the discovery nodes and executes unlearning in parallel.
-  • Decentralized: The discovery nodes publicly share Each unlearning FIM.
-Owners of each inherited node can then combine these unlearning FIMs with their own local models at will, according to specific unlearning requirements.
-This decentralized approach enables efficient and effective unlearning.
-4.4.2. Complexity
-The computational complexity of an unlearning task comprises the complexities of Steps 1, 3, and 4 (i.e., calculate unlearning FIM, merge unlearning FIMs and update models).
-The complexity of Step 2 is not considered, as the computation of the model FIM can be performed beforehand.
-The overall complexity for the unlearning task is $\mathcal{O}(|\hat{D}||L|+|\{n^{\scriptscriptstyle D}_{i}\}|+|L|)$.
-Breaking down the complexities, Eq.
-(3) and Eq.
-(4) update $|L|$ parameters per data entry, with complexities $\mathcal{O}(|\hat{D}^{\scriptscriptstyle D}_{i}||L|)$ and $\mathcal{O}(|D_{j}||L|)$, respectively, where $|\hat{D}^{\scriptscriptstyle D}_{i}|\leq|\hat{D}|$ and $|L|$ is the number of parameters in the last layer.
-The term $\mathcal{O}(|D_{j}||L|)$ is omitted as model FIMs can be calculated along with the model training prior to any unlearning tasks.
-Eq.
-(5) includes a comparison and merger of $K$ unlearning FIMs, leading to a complexity of $\mathcal{O}(K)$, where $K$ is the number of traceable discovery nodes in the unlearning graph $\mathcal{\hat{G}}$ from node $n_{j}$ and $K\leq|\{n^{\scriptscriptstyle D}_{i}\}|$.
-Eq.
-(6) updates $|L|$ model parameters, leading to a complexity of $\mathcal{O}(|L|)$.
-As the computation of the unlearning FIM for discovery nodes and the model updates can proceed in parallel, the overall complexity of the unlearning tasks is $\mathcal{O}(|\hat{D}||L|+|\{n^{\scriptscriptstyle D}_{i}\}|+|L|)$.
-5. Experimental Settings
+V-C Implementation
+System Settings. FIUn implementation varies based on system architecture:
+  • $\bullet$
+    Centralized: Model FIMs are computed and sent with the published model to the central server.
+Upon an unlearning request, the central server merges the unlearning FIMs from the discovery nodes and executes parallel unlearning.
+  • $\bullet$
+    Decentralized: Each unlearning FIM is publicly shared by discovery nodes.
+Inherited node owners can then combine these FIMs with their local models as needed, enabling efficient and effective decentralized unlearning.
+Although FIUn focuses on efficient parallel unlearning, it can be extended with auditability mechanisms to ensure verifiable unlearning in both centralized and decentralized settings.
+A practical approach is to integrate FIM commitments that record the Fisher information signatures before and after unlearning.
+Each node can compute a lightweight hash or zero-knowledge proof of its local FIM difference $\Delta F=F-\hat{F}^{D}$, which is then submitted to a verification server or blockchain ledger.
+In a centralized mode, the server validates updates by comparing FIM commitments to the expected parameter deltas from (6).
+In a decentralized mode, commitments are written on-chain, enabling independent validators to check that unlearning matches the declared forgetting requests without revealing model parameters.
+This yields an auditable trail for the “right to be forgotten” while preserving FIUn’s efficiency and scalability.
+Complexity.
+The computational complexity of an unlearning task comprises the complexities of Steps-1, 3, and 4 (i.e., calculate unlearning FIM, merge unlearning FIMs and update models, cf.
+Section V-A).
+The complexity of Step-2 is not considered, as the computation of the model FIM can be performed beforehand.
+[…]
+Equation (5) includes a comparison and merger of $K$ unlearning FIMs, leading to a complexity of $\mathcal{O}(K)$, where $K$ is the number of traceable discovery nodes in the unlearning graph $\mathcal{\hat{G}}$ from node $n_{j}$ and $K\leq|\{n^{\scriptscriptstyle D}_{i}\}|$.
+Equation (6) updates $|L|$ model parameters, leading to a complexity of $\mathcal{O}(|L|)$.
+As the computation of the unlearning FIM for discovery nodes (including the forward propagation) and the model updates can proceed in parallel, the overall complexity of the unlearning tasks is $\mathcal{O}(|\hat{D}||L_{\mathrm{all}}|+|\hat{D}||L|+|\{n^{\scriptscriptstyle D}_{i}\}|+|L|)$.
+VI Experimental Settings
 This section first presents the datasets used in the experiments, the models selected, and the performance metrics for accuracy evaluation, followed by the topology setups and the benchmarks for unlearning to evaluate the FIUn method.
-Table 1. Baseline of Federated Unlearning Performance On CIFAR-100 Unlearning Speed Comparison
-Model #$C_{f}$ Cumulative Unlearning Time (s)
-Gradient ascent Finetune Distill Re-training FIUn
+Figure 5: Models topologies considered in the experiments.
+TABLE I: Hyperparameters.
+Defaults (all models unless overridden): optimizer = SGD, learning rate = 0.01, epochs = 100, batch size = 64.
+Overrides:
+FIUn: lr = 0.1; $\tau$ = 1; $\gamma$ = 1; $\eta$ = 0.1; use layer = last.
+AlexNet: epochs = 200.
+BERT: optimizer = Adam; lr = 2e-5; epochs = 30.
+GCNN: optimizer = Adam; lr = 0.001; epochs = 30.
+ResNet18: lr = 0.1.
+TABLE II: Baseline of Various Framework Performance on CIFAR100 Unlearning Speed Comparison
+Model
+Framework
+$C_{f}$ Cumulative Unlearning Time (s)
+Gradient Ascent Fine-tuning Distill Re-training FIUn
 $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
-AlexNet 1 0.97 1.05 1.26 0.54 0.75 1.08 82.5 153.65 294.32 29.70 59.06 89.62 0.09 0.11 0.13
-2 0.99 1.17 1.39 0.57 0.77 1.15 80.49 155.63 315.52 28.69 57.81 84.97 0.11 0.16 0.13
+AlexNet
+FL 1 0.97 1.05 1.26 0.54 0.75 1.08 82.5 153.65 294.32 29.70 59.06 89.62 0.09 0.11 0.13
 10 1.1 1.27 1.68 0.53 1.95 2.31 75.3 147.6 326.8 21.96 42.37 63.82 0.11 0.14 0.15
-ResNet18 1 2.45 2.89 3.42 0.74 1.32 1.74 70.31 129.44 247.6 30.36 61.40 90.94 0.68 1.00 0.99
-2 2.02 2.2 2.78 0.76 1.22 1.79 76.55 139.55 259.53 28.30 57.39 84.06 0.69 1.03 1.09
+IL 1 1.26 1.32 1.39 1.25 2.09 3.81 108.07 212.36 424.96 22.95 43.83 63.45 0.09 0.39 0.39
+10 1.29 1.39 1.54 1.26 2.14 3.24 109.3 214.35 421.85 17.35 38.32 57.94 0.12 0.28 0.29
+DDPL 1 1.42 1.47 1.56 1.53 1.83 1.79 326.64 108.13 121.98 30.13 31.26 31.14 0.23 0.14 0.09
+10 1.56 1.63 1.75 1.74 2.03 2.01 114.32 111.3 111.19 21.03 22.96 22.10 0.24 0.14 0.10
+TL 1 1.14 1.21 1.3 1.16 2.19 1.98 112.87 220.13 226.31 20.31 40.49 42.36 0.09 0.11 0.11
+10 1.26 1.34 1.36 1.25 2.06 2.01 109.44 217.32 223.91 19.42 39.46 39.35 0.08 0.13 0.11
+ResNet18
+FL 1 2.45 2.89 3.42 0.74 1.32 1.74 70.31 129.44 247.6 30.36 61.40 90.94 0.68 1.00 0.99
 10 2.02 2.58 3.67 0.85 1.54 1.84 75.12 150.35 267.34 22.03 44.83 66.30 0.76 1.14 1.04
-(a) FL in ResNet18 of #$C_{f}=1$
-(b) FL in ResNet18 of #$C_{f}=10$
-(c) FL in AlexNet of #$C_{f}=1$
-(d) FL in AlexNet of #$C_{f}=10$
-Figure 5. Baseline of Federated Unlearning Performance On CIFAR-100.
-5.1. Datasets and Models
-Datasets. The experiment uses two datasets to evaluate the proposed FIUn method. These two datasets are benchmarks for image classification tasks, encompassing diverse image classification tasks.
-  • CIFAR100 (Krizhevsky et al., 2009).
-This dataset consists of 60,000 color images in 100 classes, with 600 images per class.
-The images are of size $32\times 32$ pixels.
-The dataset is divided into 50,000 training images and 10,000 test images.
-Each class in the CIFAR-100 dataset belongs to one of 20 superclasses.
-There is a coarse label (superclass) and a fine label (specific class) per image.
-  • TinyImageNet (Le and Yang, 2015).
-This dataset contains 200 classes, each with 500 training images, 50 validation images, and 50 test images, with a total of up to 100,000 images.
-The images are resized to $64\times 64$ pixels, making them smaller than those in the original ImageNet dataset but larger than those in CIFAR-100.
+IL 1 1.52 1.63 1.84 1.35 2.65 3.86 98.51 170.66 317.53 26.35 47.12 66.93 0.30 0.80 0.85
+10 1.33 1.45 1.62 1.47 2.24 3.74 81.74 143.53 275.34 18.93 39.93 58.93 0.32 0.88 0.89
+DDPL 1 1.18 1.25 1.34 1.57 1.43 1.51 72.34 71.87 71.39 32.53 31.95 32.18 0.64 0.34 0.30
+10 1.34 1.42 1.51 2.04 1.85 1.97 71.41 66.56 70.41 22.60 21.92 24.19 0.68 0.34 0.34
+TL 1 1.39 1.48 1.52 1.37 3.67 3.25 95.4 163.54 168.51 22.21 41.52 42.93 0.15 0.24 0.24
+10 1.35 1.51 1.46 1.5 5.53 5.54 81.15 145.73 151.52 19.10 42.60 41.36 0.16 0.27 0.27
+TABLE III: Transfer Unlearning Performance on CIFAR100
+Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
+Re-training FIUn
+$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
+AlexNet
+1 $AD_{r}\uparrow$ 97.16 95.61 92.20 98.17 95.98 91.32 91.40 86.24 83.99 20.31 40.49 42.36
+${AD}_{f}\downarrow$ 96.66 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.09 0.11 0.11
+2 $AD_{r}\uparrow$ 97.16 95.61 92.20 96.88 96.68 93.16 82.36 80.47 79.99 21.20 40.56 41.92
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.08 0.13 0.11
+4 $AD_{r}\uparrow$ 97.16 95.61 92.20 94.67 96.45 93.45 71.59 71.43 76.06 20.30 39.76 40.60
+${AD}_{f}\downarrow$ 96.61 98.30 98.30 0.00 0.00 0.00 0.00 0.00 0.00 0.09 0.11 0.11
+ResNet18
+1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 89.09 93.51 95.22 22.21 41.52 42.93
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.15 0.24 0.24
+2 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 85.89 94.89 95.79 21.19 41.72 43.20
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.17 0.23 0.26
+4 $AD_{r}\uparrow$ 99.96 99.99 99.99 99.99 99.99 99.99 80.20 94.82 97.03 22.03 42.26 43.25
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.17 0.25 0.25
+VI-A Datasets and Models
+Datasets. The experiment uses three datasets to evaluate the proposed FIUn method. These three datasets are benchmarks for image classification tasks, encompassing diverse image classification tasks.
+  • $\bullet$
+    CIFAR100.
+This dataset includes 60,000 color images across 100 classes, each with 600 images, sized at $32\times 32$ pixels.
+It comprises 50,000 training and 10,000 test images.
+Each class belongs to one of 20 superclasses, with both a coarse (superclass) and fine (specific class) label per image.
+  • $\bullet$
+    TinyImageNet.
+This dataset includes 200 classes, with 500 training images, 50 validation images, and 50 test images per class, totaling 100,000 images.
+The images are resized to $64\times 64$ pixels, larger than CIFAR100 but smaller than the original ImageNet.
+  • $\bullet$
+    Yahoo!
+Answers.
+his text dataset is a large public Q&A dataset widely used in natural language processing (NLP) research for tasks like text classification, question answering, and sentiment analysis.
+It contains categorized question-answer pairs from Yahoo!
+Answers.
 Models.
-We use the AlexNet model (Iandola et al., 2016), ResNet18 model (Odusami et al., 2021) and DenseNet161 model (Talo, 2019) to evaluate the impact of our proposed FIUn method on the accuracy of the unlearned label set $C_{f}$ and the retained label set $C_{r}$.
+We use the AlexNet, ResNet18, DenseNet161, BERT-base-cased language model and GCNN to evaluate the impact of our proposed FIUn method on the accuracy of the unlearned label set $C_{f}$ and the retained label set $C_{r}$.
 Performance Metrics.
 We use the training datasets to check the accuracy of the model to evaluate its practicality in experiments.
-This is a common and reasonable approach since it directly evaluates if the removed information still influences the model, as widely accepted in the field (Tarun et al., 2023; Kim et al., 2024; Golatkar et al., 2020).
+This is a common and reasonable approach since it directly evaluates if the removed information still influences the model, as widely accepted in the field [tarun2023fast, golatkar2020eternal].
+  • $\bullet$
     Accuracy on unlearned labels ($AD_{f}$).
-The accuracy of the unlearned label set in the unlearning model ideally would be close to zero.
-We use the training datasets to test the effectiveness of unlearning.
-  • Accuracy on retained labels ($AD_{r}$). The accuracy of a retained label set of the unlearning model. It is expected to be close to the accuracy of the original model before unlearning.
+The accuracy of the unlearned label set in the unlearning model should ideally approach 0%.
+This represents a scenario where the softmax layer assigns negligible probabilities to unlearned classes.
+When applying the Top-$N$ strategy, the likelihood of selecting unlearned classes becomes extremely small, effectively resulting in 0% accuracy.
+We evaluate this by testing the unlearning model on training datasets to assess the effectiveness of the unlearning process.
+  • $\bullet$
+    Accuracy on retained labels ($AD_{r}$). The accuracy of the retained label set should closely match the original model’s accuracy before unlearning.
+  • $\bullet$
     Cumulative unlearning time. The cumulative time required for each model to unlearn labels during the training process in machine learning and deep learning.
+  • $\bullet$
     Accuracy difference index ($\bigtriangleup_{acc}$).
-The difference between $AD_{r}$ and $AD_{f}$ also serves as a direct metric to evaluate unlearning performance.
-The greater the difference, the more effective the unlearning is, with a maximum 1.
-Table 2. Federated Unlearning Performance On TinyImageNet
-Model #$C_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
+The difference between $AD_{r}$ and $AD_{f}$ directly measures unlearning effectiveness, with a larger difference indicating more effective unlearning, up to a maximum value of 1.
+(a) FL in ResNet18 of $C_{f}=1$
+(b) FL in ResNet18 of $C_{f}=10$
+(c) FL in AlexNet of $C_{f}=1$
+(d) FL in AlexNet of $C_{f}=10$
+(e) DDPL in ResNet18 of $C_{f}=1$
+(f) DDPL in ResNet18 of $C_{f}=10$
+(g) DDPL in AlexNet of $C_{f}=1$
+(h) DDPL in AlexNet of $C_{f}=10$
+Figure 6: Baseline Unlearning Accuracy of FL and DDPL Frameworks on CIFAR100.
+(a) IL in ResNet18 of $C_{f}=1$
+(b) IL in ResNet18 of $C_{f}=10$
+(c) IL in AlexNet of $C_{f}=1$
+(d) IL in AlexNet of $C_{f}=10$
+(e) TL in ResNet18 of $C_{f}=1$
+(f) TL in ResNet18 of $C_{f}=10$
+(g) TL in AlexNet of $C_{f}=1$
+(h) TL in AlexNet of $C_{f}=10$
+Figure 7: Baseline Unlearning Accuracy of IL and TL Frameworks on CIFAR100.
+VI-B Framework Setup and Benchmarks
+Targeted Frameworks. We detail setups for four learning frameworks with DAG, all adopting the structure shown in Fig. 5. These frameworks vary in their dataset and label distribution conditions.
+  • $\bullet$
+    Federated Unlearning.
+(Yellow in Fig.
+5) The CIFAR100 and TinyImageNet datasets are each randomly divided into five parts.
+When the model $w_{g}$ contains labels that need to be unlearned, the models $w_{a}$ and $w_{b}$ that inherit from it need to perform the unlearning operation.
+  • $\bullet$
+    Distributed Data-Parallel Unlearning.
+(Red in Fig.
+5) The CIFAR100 and TinyImageNet datasets are each randomly divided into three parts, each executing a sub-task in parallel.
+When the model $w_{g}$ contains labels that need to be unlearned, sub-tasks $w_{a}$ and $w_{b}$ perform the unlearning operation in parallel.
+  • $\bullet$
+    Incremental Unlearning.
+(Blue in Fig.
+5) In the CIFAR100 and TinyImageNet datasets, the initial models are trained on labels 0-90 and 0-190, respectively, with each inheriting model adding two more classes.
+Models $w_{g}$, $w_{a}$, and $w_{b}$ are selected for experimental demonstration.
+  • $\bullet$
+    Transfer Unlearning.
+(Red in Fig.
+5) Since TL involves fine-tuning the last layer, adjustments are focused on this layer.
+In experiments, a model $w_{g}$ is trained on CIFAR100 labels 0-90, then transferred to models $w_{a}$ and $w_{b}$ trained on labels 0-92 and 0-98, respectively.
+Similarly, in TinyImageNet, model $w_{g}$ is trained on labels 0-190 and transferred to models $w_{a}$ and $w_{b}$ trained on labels 0-192 and 0-198.
+TABLE IV: Distributed Data-Parallel Unlearning Performance on CIFAR100
+Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
 Re-training FIUn
 $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
-DenseNet161 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.81 99.99 99.99 98.80 88.03 77.64 76.02 999.41 1996.43 2997.64 2.51 4.59 4.95
-${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00
-4 ${AD}_{r}\uparrow$ 99.99 99.99 99.80 99.99 99.99 99.51 83.01 71.43 71.00 1004.64 2006.31 3006.14 3.16 4.69 4.89
-${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00
-6 ${AD}_{r}\uparrow$ 99.99 99.99 99.96 99.99 99.99 99.16 80.00 67.67 66.58 1019.43 2020.16 3018.64 3.10 4.96 4.75
-${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00
-10 ${AD}_{r}\uparrow$ 99.99 99.99 99.72 99.71 99.99 99.99 71.68 58.01 55.80 1031.43 2029.64 3030.64 2.95 4.85 4.89
-${AD}_{f}\downarrow$ 99.99 99.99 98.82 0.00 0.00 0.00 0.00 0.00 0.00
-ResNet18 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.73 99.97 99.46 112.95 216.39 314.29 1.35 3.68 3.42
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-4 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.17 99.98 99.98 109.89 209.31 310.42 1.63 3.26 3.57
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-6 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 99.54 99.98 99.98 107.39 208.12 305.32 1.46 3.37 3.51
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-10 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.98 99.99 99.99 104.21 201.36 302.74 1.52 3.21 3.36
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-Table 3. Incremental Unlearning Performance On TinyImageNet
-Model #$C_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
-Re-training FIUn
-$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
-DenseNet161 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.81 99.99 99.99 98.80 89.76 78.50 73.87 991.43 1995.36 3989.47 2.75 4.85 4.43
-${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00
-4 ${AD}_{r}\uparrow$ 99.99 99.99 99.80 99.99 99.99 99.51 83.98 65.87 59.11 1006.32 2004.75 4009.41 2.64 4.43 4.64
-${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00
-6 ${AD}_{r}\uparrow$ 99.99 99.99 99.96 99.99 99.99 99.16 86.81 61.02 55.04 1019.43 2019.50 4019.46 2.36 4.47 4.74
-${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00
-10 ${AD}_{r}\uparrow$ 99.99 99.99 99.72 99.99 99.99 99.99 85.95 54.61 45.46 1035.36 2041.78 4039.75 2.65 4.46 4.75
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-ResNet18 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 98.66 96.33 96.07 120.32 221.75 439.85 1.37 3.75 3.43
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-4 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 98.84 96.34 96.70 115.32 217.33 435.17 1.41 3.34 3.64
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-6 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 98.82 99.99 98.59 110.32 211.35 431.37 1.47 3.46 3.75
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-10 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.33 97.51 98.45 108.32 207.43 428.43 1.32 3.56 3.57
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-5.2. Learning Framework Setup and Benchmarks
-Figure 6. Models topologies considered in the experiments.
-Targeted Frameworks. We detailed setups for four learning frameworks under the UMIG with varying dataset and label distribution conditions.
-  • Federated Unlearning.
-This process uses the structure shown in Figure 6.
-The CIFAR100 and TinyImageNet datasets are each randomly divided into five parts.
-When the model $w_{g}$ contains labels that need to be unlearned, the models that inherit from it need to perform the unlearning operation.
-We select models $w_{g}$, $w_{a}$, and $w_{b}$ for experimental demonstration.
-  • Distributed Data-Parallel Unlearning.
-This process uses the structure shown in Figure 6.
-The CIFAR100 and TinyImageNet datasets are each randomly divided into three parts, each executing a sub-task in parallel.
-When the model $w_{g}$ contains labels that need to be unlearned, each sub-task performs the unlearning operation in parallel.
-We select models $w_{g}$, $w_{a}$, and $w_{b}$ for experimental demonstration.
-  • Incremental Unlearning.
-This process uses the structure shown in Figure 6.
-In the CIFAR100 dataset, the first model is trained on data with labels 0-90, and each inheriting model’s training dataset progressively includes two additional classes.
-In the TinyImageNet dataset, the first model is trained on data with labels 0-190, and similarly, each inheriting model’s training dataset progressively includes two additional classes.
-We select models $w_{g}$, $w_{a}$, and $w_{b}$ for experimental demonstration.
-  • Transfer Unlearning.
-Since TL involves fine-tuning the last layer parameters, adjustments are focused on the last layer of the model.
-Experiments, as illustrated in Figure 6, involve the CIFAR100 dataset where a model $w_{g}$ is trained on labels 0-90 and then transferred to models $w_{a}$ and $w_{b}$ trained on labels 0-92 and 0-98, respectively.
-Similarly, for the TinyImageNet dataset, model $w_{g}$ is trained on labels 0-190 and then transferred to models $w_{a}$ and $w_{b}$, which are trained on labels 0-192 and 0-198, respectively.
+AlexNet
+1 $AD_{r}\uparrow$ 98.01 99.96 99.96 99.96 99.96 99.96 97.23 98.85 97.66 30.13 31.26 31.14
+${AD}_{f}\downarrow$ 99.99 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00 0.23 0.14 0.09
+2 $AD_{r}\uparrow$ 99.96 99.96 99.96 99.96 99.96 99.96 96.14 97.07 96.71 26.43 26.47 26.41
+${AD}_{f}\downarrow$ 99.96 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00 0.27 0.16 0.11
+4 $AD_{r}\uparrow$ 98.01 99.96 99.96 99.96 99.96 99.96 94.58 90.41 90.97 24.33 24.36 24.37
+${AD}_{f}\downarrow$ 99.09 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00 0.23 0.14 0.10
+ResNet18
+1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 91.87 94.67 93.67 32.53 31.95 32.18
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.64 0.34 0.30
+2 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 89.94 94.58 92.40 28.28 27.22 28.47
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.66 0.34 0.30
+4 $AD_{r}\uparrow$ 99.96 89.20 88.52 99.96 99.96 86.44 93.95 89.26 91.77 25.37 24.97 25.13
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.66 0.35 0.31
 Benchmark.
@@
 We remove the data with unlearned labels from each model’s dataset.
-The remaining data is then used to individually train each model.
+The remaining data is then used to train each model individually.
 The models are trained sequentially according to their inheritance relationships.
-Hyperparameters.
-The hyperparameters used in the experiment are shown in Appendix .1.
-The table lists the learning rates, optimizers, and other hyperparameters for different models, as well as the hyperparameters of our proposed FIUn method.
-Throughout the experimental process, all the hyperparameters remain unchanged to ensure a fair comparison of our method among different models.
-Table 4. Transfer Unlearning Performance On CIFAR-100
-Model #$C_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
+Hyperparameters. Table I summarizes the hyperparameters for various models and our FIUn method. All hyperparameters remain fixed during experiments to ensure fair comparisons across models.
+VI-C Label Category Settings
+The unlearned label categories, denoted as $C_{f}$, may vary across different models and data distributions.
+For instance, model $w_{a}$ aims to unlearn labels 2, 3, and 4, while model $w_{b}$ intends to unlearn labels 2, 3, and 6.
+There is an overlap between the unlearned label categories of model $w_{a}$ and model $w_{b}$, specifically labels 2 and 3.
+We refer to the degree of overlapping among these unlearned label categories as $\cap_{f}$.
+The unlearning process is applied to individual and combined labels.
+For CIFAR100 and Yahoo!
+Answers, we select combinations of 1, 2, and 4 categories, while for TinyImageNet, we select 1, 4, and 6 categories.
+The overlap of unlearned labels is divided into 60%, 40%, and 20%.
+VII Evaluation
+We use an NVIDIA 4090 GPU to evaluate FIUn’s impact on the accuracy of unlearned and retained labels across four frameworks: FL, DDPL, IL, and TL.
+Accuracy and time consumption during unlearning are assessed by comparing FIUn with existing methods, followed by an analysis of re-training, the most effective (but inefficient) unlearning benchmark, focusing on inheritance depth and parameter layers impact on unlearning effectiveness.
+VII-A General Label Unlearning Analysis
+We consider existing emerging unlearning methods for comparison: 1) fine-tuning, which adjusts the original model using the remaining dataset [odusami2021analysis], 2) gradient ascent (GA), employing negative gradients for unlearning [ga], and 3) distill, which distills knowledge from the original model into a student model using the remaining data [distil1Unl].
+Table II reports unlearning speed and Fig.6 shows unlearning accuracy $\bigtriangleup_{acc}$; Fig.7 compares $\bigtriangleup_{acc}$ under IL and TL.
+Our method improves unlearning speed by up to 99% over alternatives and maintains a consistent accuracy advantage: for single-label and up to 10 classes, GA and fine-tuning finish in about 2 seconds, distillation and retraining are slower, while our approach averages under 1 second.
+In accuracy, fine-tuning and GA remain suboptimal, and our method matches or surpasses distillation across most architectures (with a slight dip on AlexNet).
+Overall, it delivers the fastest cumulative unlearning across IMNs.
+Building on this, we next compare with retraining, which is widely regarded as the gold standard for unlearning and retention accuracy [sisa]; see Tables III and VI, with additional results in Appendix A.
+TABLE V: Federated Unlearning Performance on TinyImageNet
+Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
 Re-training FIUn
 $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
-AlexNet 1 ${AD}_{r}\uparrow$ 97.16 95.61 92.20 98.17 95.98 91.32 91.40 86.24 83.99 20.31 40.49 42.36 0.09 0.11 0.11
-${AD}_{f}\downarrow$ 96.66 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-2 ${AD}_{r}\uparrow$ 97.16 95.61 92.20 96.88 96.68 93.16 82.36 80.47 79.99 21.20 40.56 41.92 0.08 0.13 0.11
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-4 ${AD}_{r}\uparrow$ 97.16 95.61 92.20 94.67 96.45 93.45 71.59 71.43 76.06 20.30 39.76 40.60 0.09 0.11 0.11
-${AD}_{f}\downarrow$ 96.61 98.30 98.30 0.00 0.00 0.00 0.00 0.00 0.00
-10 ${AD}_{r}\uparrow$ 97.16 95.61 92.20 99.99 99.99 99.99 75.97 69.56 76.15 19.42 39.46 39.35 0.08 0.13 0.11
-${AD}_{f}\downarrow$ 98.57 99.28 99.28 0.00 0.00 0.00 20.28 20.28 20.28
-ResNet18 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 89.09 93.51 95.22 22.21 41.52 42.93 0.15 0.24 0.24
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-2 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 85.89 94.89 95.79 21.19 41.72 43.20 0.17 0.23 0.26
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-4 ${AD}_{r}\uparrow$ 99.96 99.99 99.99 99.99 99.99 99.99 80.20 94.82 97.03 22.03 42.26 43.25 0.17 0.25 0.25
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-10 ${AD}_{r}\uparrow$ 99.96 99.99 99.99 99.99 99.99 99.99 75.36 98.07 98.96 19.10 42.60 41.36 0.16 0.27 0.27
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-Table 5. Distributed Data-Parallel Unlearning Performance On CIFAR-100
-Model #$C_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
+DenseNet161
+1 $AD_{r}\uparrow$ 99.99 99.99 99.81 99.99 99.99 98.80 88.03 77.64 76.02 999.41 1996.43 2997.64
+${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00 2.51 4.59 4.95
+4 $AD_{r}\uparrow$ 99.99 99.99 99.80 99.99 99.99 99.51 83.01 71.43 71 1004.64 2006.31 3006.14
+${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00 3.16 4.69 4.89
+6 $AD_{r}\uparrow$ 99.99 99.99 99.96 99.99 99.99 99.16 80.00 67.67 66.58 1019.43 2020.16 3018.64
+${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00 3.10 4.96 4.75
+ResNet18
+1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.73 99.97 99.46 112.95 216.39 314.29
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 1.35 3.68 3.42
+4 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.17 99.98 99.98 109.89 209.31 310.42
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 1.63 3.26 3.57
+6 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 99.54 99.98 99.98 107.39 208.12 305.32
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 1.46 3.37 3.51
+TABLE VI: Incremental Unlearning Performance on TinyImageNet
+Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
 Re-training FIUn
 $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
-AlexNet 1 ${AD}_{r}\uparrow$ 98.01 99.96 99.96 99.96 99.96 99.96 97.23 98.85 97.66 30.13 31.26 31.14 0.09 0.14 0.23
-${AD}_{f}\downarrow$ 99.99 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00
-2 ${AD}_{r}\uparrow$ 99.96 99.96 99.96 99.96 99.96 99.96 96.14 97.07 96.71 26.43 26.47 26.41 0.11 0.16 0.27
-${AD}_{f}\downarrow$ 99.96 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00
-4 ${AD}_{r}\uparrow$ 98.01 99.96 99.96 99.96 99.96 99.96 94.58 90.41 90.97 24.33 24.36 24.37 0.10 0.14 0.23
-${AD}_{f}\downarrow$ 99.09 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00
-10 ${AD}_{r}\uparrow$ 98.01 99.96 99.96 99.96 99.96 99.96 91.53 76.19 80.70 21.03 22.96 22.10 0.10 0.14 0.24
-${AD}_{f}\downarrow$ 97.15 99.96 99.96 0.00 0.00 0.00 9.73 0.00 11.08
-ResNet18 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 91.87 94.67 93.67 32.53 31.95 32.18 0.30 0.34 0.64
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-2 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 89.94 94.58 92.40 28.28 27.22 28.47 0.30 0.34 0.66
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-4 ${AD}_{r}\uparrow$ 99.96 89.20 88.52 99.96 99.96 86.44 93.95 89.26 91.77 25.37 24.97 25.13 0.31 0.35 0.66
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-10 ${AD}_{r}\uparrow$ 99.96 89.20 88.52 99.96 99.96 99.96 90.44 95.79 94.19 22.60 21.92 24.19 0.34 0.34 0.68
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 9.73 0.00 0.00
-Table 6. Federated Unlearning Performance On TinyImageNet With Multiple Label Distribution
-Model #flimit-$\#\cap_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
-Re-training FIUn
-$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
-DenseNet161 60% $AD_{r}\uparrow$ 99.99 99.99 99.81 98.80 99.99 99.99 96.39 84.75 77.89 3000.66 1998.46 994.32 4.69 2.39 2.30
-${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00
-40% $AD_{r}\uparrow$ 99.50 99.99 99.80 99.51 99.99 99.99 96.64 85.40 85.00 3009.53 1010.45 2011.14 5.12 2.69 2.14
-${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00
-20% $AD_{r}\uparrow$ 99.99 99.99 99.96 99.16 99.99 99.99 96.22 81.49 78.83 3022.64 2020.67 1021.34 4.98 2.64 2.59
-${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00
-0% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.71 96.01 83.04 79.15 3039.64 2041.32 1046.14 4.96 2.14 2.50
-${AD}_{f}\downarrow$ 99.99 99.99 98.82 0.00 0.00 0.00 0.00 0.00 0.00
-ResNet18 60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 95.23 98.54 98.44 349.32 246.43 123.74 3.16 1.65 1.68
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-40% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 96.98 99.44 99.49 355.32 247.64 128.43 2.97 1.46 1.59
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 94.43 98.86 99.38 354.83 246.43 129.35 2.89 1.54 1.38
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-0% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 97.36 99.34 99.32 351.38 244.34 125.32 3.10 1.69 1.39
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-Table 7. Distributed Data-Parallel Unlearning Performance On CIFAR-100 With Multiple Label Distribution
-Model #flimit-$\#\cap_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
-Re-training FIUn
-$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
-DenseNet161 60% $AD_{r}\uparrow$ 99.99 99.99 99.81 98.80 99.99 99.99 95.18 97.07 99.99 999.41 996.43 997.64 0.15 0.08 0.09
-${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00
-40% $AD_{r}\uparrow$ 99.99 99.99 99.80 99.51 99.99 99.99 90.53 97.07 98.82 1004.64 1006.31 1006.14 0.19 0.09 0.11
-${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00
-20% $AD_{r}\uparrow$ 99.99 99.99 99.96 99.16 99.99 99.99 95.29 97.07 99.89 1019.43 1020.16 1018.64 0.18 0.07 0.10
-${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00
-0% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 87.15 97.07 96.36 1031.43 1029.64 1030.64 0.18 0.10 0.11
-${AD}_{f}\downarrow$ 99.99 99.99 98.82 0.00 0.00 0.00 0.00 0.00 0.00
-ResNet18 60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 97.62 88.94 87.87 30.36 33.18 37.31 0.60 0.32 0.29
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-40% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 96.50 89.35 87.89 28.30 29.13 52.33 0.59 0.34 0.30
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 95.31 88.17 93.06 26.39 23.37 56.61 0.58 0.28 0.30
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-0% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 88.87 88.86 89.16 22.03 24.36 46.42 0.64 0.37 0.32
-${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
-5.3. Label Category Settings
-The unlearned label categories are represented as #$\#C_{f}$.
-The distribution of labels among different models may also vary in different data distributions.
-For instance, model $w_{a}$ aims to unlearn labels 1, 2, 3, and 4, while model $w_{b}$ intends to unlearn labels 2, 3, 5, and 6.
-In this case, there is an overlap between the unlearned label categories of model $w_{a}$ and model $w_{b}$, specifically labels 2 and 3.
-We refer to the degree of overlapping among these unlearned label categories as #flimit-$\#\cap_{f}$.
-The unlearning process is applied to individual categories and various combinations of labels.
-For the CIFAR-100 dataset, the combinations we select include 1, 2, 4, and 10 categories.
-For the TinyImageNet dataset, the selected combinations include 1, 4, 6, and 10 categories.
-The overlap of the unlearned label categories is divided into different proportions, including 60%, 40%, 20%, and 0%.
-6. Evaluation
-In this section, we use NVIDIA 4090 GPU to set up a test platform that demonstrates the impact of the FIUn method on the accuracy of unlearned and retained labels in four frameworks: FL, DDPL, IL and TL.
-When assessing accuracy and time consumption across these learning frameworks during the model unlearning process, we compare our proposed FIUn method with existing unlearning methods and use re-training as a benchmark.
-We also analyze the impact of the inheritance depth and parameter layers of the models on unlearning effectiveness.
+DenseNet161
+1 $AD_{r}\uparrow$ 99.99 99.99 99.81 99.99 99.99 98.80 89.76 78.50 73.87 991.43 1995.36 3989.47
+${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00 2.75 4.85 4.43
+4 $AD_{r}\uparrow$ 99.99 99.99 99.80 99.99 99.99 99.51 83.98 65.87 59.11 1006.32 2004.75 4009.41
+${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00 2.64 4.43 4.64
+6 $AD_{r}\uparrow$ 99.99 99.99 99.96 99.99 99.99 99.16 86.81 61.02 55.04 1019.43 2019.50 4019.46
+${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00 2.36 4.47 4.74
+ResNet18
+1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 98.66 96.33 96.07 120.32 221.75 439.85
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 1.37 3.75 3.43
+4 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 98.84 96.34 96.70 115.32 217.33 435.17
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 1.41 3.34 3.64
+6 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 98.82 99.99 98.59 110.32 211.35 431.37
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 1.47 3.46 3.75
 (a) FL in CIFAR100 ResNet18
@@
 (d) FL TinyImageNet by DenseNet161
-Figure 7. Overlap #$\cap_{f}$ analysis, #$C_{f}$=15.
-6.1. General Label Unlearning Analysis
-We consider existing emerging unlearning methods for comparison: 1) Fine-tuning, which adjusts the original model using the remaining dataset (Odusami et al., 2021), 2) Gradient ascent (GA), employing negative gradients for unlearning (Zhang et al., 2024), and 3) Distill, which distills knowledge from the original model into a student model using the remaining data (Sinha et al., 2023).
-Performance comparisons are depicted in Figure 5 for unlearning accuracy $\bigtriangleup_{acc}$ and detailed in Table 1 for unlearning speed.
-The results demonstrate that our method not only enhances unlearning speed by up to 99% over alternative unlearning methods but also consistently outperforms them in unlearning accuracy.
-These results indicate that our method exhibits outstanding effectiveness and efficiency in handling the complexity of model unlearning.
-In what follows, we evaluate the proposed parallel unlearning method in comparison with the most robust method—re-training.
-Within all considered learning frameworks, our proposed FIUn method achieves a maximum increase in unlearning speed of 99% compared to the re-training benchmark across different datasets and models.
-The detailed results are summarized in Tables 2 to 5, with more experiments provided in Appendix .2.1.
+Figure 8: Overlap $\cap_{f}$ analysis, $C_{f}$=15. MFIM stands for Merging FIM function.
 1) Single-label unlearning.
@@
 For all frameworks and model types in TinyImageNet, the $AD_{f}$ metric also reaches 0, and the average $AD_{r}$ metric reaches 79.49%.
-This significantly validates the effectiveness of our method.
+This validates the effectiveness of our method.
 2) Multi-label unlearning.
-For the CIFAR100 dataset, with #$C_{f}=2,4,10$, the average $AD_{f}$ metric across all frameworks and model types is 1.93%, while the average $AD_{r}$ metric reaches 86.01%.
+For the CIFAR100 dataset, with $C_{f}=2,4$, the average $AD_{f}$ metric across all frameworks and model types is 1.93%, while the average $AD_{r}$ metric reaches 86.01%.
 For the TinyImageNet dataset, the average $AD_{f}$ metric is 0.21%, and the average $AD_{r}$ metric is 83.54%.
@@
 However, the $AD_{r}$ performance of the AlexNet and DenseNet161 models shows a downward trend, while the $AD_{r}$ performance of the ResNet18 model remains stable.
+TABLE VII: Distributed Data-Parallel Unlearning Performance on CIFAR100 with Multiple Label Distribution
+Model $\cap_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
+Re-training FIUn
+$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
+DenseNet161
+60% $AD_{r}\uparrow$ 99.99 99.99 99.81 98.80 99.99 99.99 95.18 97.07 99.99 999.41 996.43 997.64
+$AD_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00 0.15 0.08 0.09
+40% $AD_{r}\uparrow$ 99.99 99.99 99.80 99.51 99.99 99.99 90.53 97.07 98.82 1004.64 1006.31 1006.14
+$AD_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00 0.19 0.09 0.11
+20% $AD_{r}\uparrow$ 99.99 99.99 99.96 99.16 99.99 99.99 95.29 97.07 99.89 1019.43 1020.16 1018.64
+$AD_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00 0.18 0.07 0.10
+ResNet18
+60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 97.62 88.94 87.87 30.36 33.18 37.31
+$AD_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.60 0.32 0.29
+40% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 96.50 89.35 87.89 28.30 29.13 52.33
+$AD_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.59 0.34 0.30
+20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 95.31 88.17 93.06 26.39 23.37 56.61
+$AD_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.58 0.28 0.30
+TABLE VIII: Federated Unlearning Performance on TinyImageNet with Multiple Label Distribution
+Model $\cap_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
+Re-training FIUn
+$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
+DenseNet161
+60% $AD_{r}\uparrow$ 99.99 99.99 99.81 98.80 99.99 99.99 96.39 84.75 77.89 3000.66 1998.46 994.32
+${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00 4.69 2.39 2.30
+40% $AD_{r}\uparrow$ 99.50 99.99 99.80 99.51 99.99 99.99 96.64 85.40 85.00 3009.53 1010.45 2011.14
+${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00 5.12 2.69 2.14
+20% $AD_{r}\uparrow$ 99.99 99.99 99.96 99.16 99.99 99.99 96.22 81.49 78.83 3022.64 2020.67 1021.34
+${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00 4.98 2.64 2.59
+ResNet18
+60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 95.23 98.54 98.44 349.32 246.43 123.74
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 3.16 1.65 1.68
+40% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 96.98 99.44 99.49 355.32 247.64 128.43
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 2.97 1.46 1.59
+20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 94.43 98.86 99.38 354.83 246.43 129.35
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 2.89 1.54 1.38
 Further observation reveals that the number of parameters in the last layer of DenseNet161 is 4.3 times that of ResNet18, and the number of parameters in the last layer of AlexNet is 7.9 times that of ResNet18.
-These results indicate that as the number of model parameters and the classes of forgotten labels increases, the $AD_{r}$ performance may decline.
-This further suggests that while our proposed method effectively unlearns specified labels, it can somewhat affect the accuracy of retained labels.
-(a) FL TinyImageNet by ResNet18
-(b) FL CIFAR100 by ResNet18
-(c) FL TinyImageNet by DenseNet161
-(d) FL CIFAR100 by DenseNet161
-(e) IL TinyImageNet by ResNet18
-(f) IL CIFAR100 by ResNet18
-(g) IL TinyImageNet by DenseNet161
-(h) IL CIFAR100 by DenseNet161
-Figure 8. Inheritance depth analysis #$C_{f}=4$.
-(a) FL AlexNet in CIFAR100
-(b) FL CIFAR100 by ResNet18
-(c) IL CIFAR100 by ResNet18
-(d) IL CIFAR100 by AlexNet
-(e) FL TinyImageNet by DenseNet161
-(f) FL TinyImageNet by ResNet18
-(g) IL TinyImageNet by ResNet
-(h) IL TinyImageNet by DenseNet161
-Figure 9. Unlearning layer analysis #$C_{f}$=15.
-6.2. Merged Label Unlearning Analysis
-This section assesses the applicability of the merging FIM function to models with various distributions of unlearned labels.
-We analyze the unlearning performance of different types of models across various datasets within the FL and DDPL frameworks.
-As demonstrated in Tables 6 and 7, with more experiments provided in Appendix .2.2, the average $AD_{r}$ is 89.72%, and the average $AD_{f}$ is 5.62%, confirming the unlearning effectiveness of our method.
-We further validate that in the FIM, the impact of data on model parameters remains relatively stable.
-We also analyze whether our proposed merging FIM function can handle the merged label unlearning task within the FL framework.
-As illustrated in Figure 7, in the AlexNet and DenseNet161 models, our method shows an average performance improvement of 88.54%, compared to using FIM directly.
-In the ResNet18 model, our method’s performance closely matches that of FIM.
-Additionally, our proposed merged label unlearning is, on average, 2.3 times faster than general label unlearning.
-This further validates the effiency of our method.
-6.3. Inherited Depth of Model Analysis
-We further examine whether the performance of FIUn is influenced by the increasing depth of inheritance within the FL and IL frameworks. All experiments are conducted with the same hyperparameters.
+These results indicate that as the number of model parameters and the classes of unlearned labels increases, the $AD_{r}$ performance may decline.
+Takeaway–faster and better.
+FIUn outperforms others in unlearning speed by 99%, averaging under one second per model, while effectively removing targeted labels and preserving retained accuracy, matching the effectiveness of the re-training method, which is considered the most effective.
+(a) FL, TinyImageNet, ResNet18
+(b) FL, CIFAR100, ResNet18
+(c) FL, TinyImageNet, DenseNet161
+(d) FL, CIFAR100, DenseNet161
+(e) IL, TinyImageNet, ResNet18
+(f) IL, CIFAR100, ResNet18
+(g) IL, TinyImageNet, DenseNet161
+(h) IL, CIFAR100, DenseNet161
+Figure 9: Inheritance depth analysis $C_{f}=1$.
+(a) FL, TinyImageNet, ResNet18
+(b) FL, CIFAR100, ResNet18
+(c) FL, TinyImageNet, DenseNet161
+(d) FL, CIFAR100, DenseNet161
+(e) IL, TinyImageNet, ResNet18
+(f) IL, CIFAR100, ResNet18
+(g) IL, TinyImageNet, DenseNet161
+(h) IL, CIFAR100, DenseNet161
+Figure 10: Inheritance depth analysis $C_{f}=4$.
+(a) FL, AlexNet, CIFAR100
+(b) FL, CIFAR100, ResNet18
+(c) IL, CIFAR100, ResNet18
+(d) IL,CIFAR100, AlexNet
+(e) FL, TinyImageNet, DenseNet161
+(f) FL, TinyImageNet, ResNet18
+(g) IL, TinyImageNet, ResNet18
+(h) IL, TinyImageNet, DenseNet161
+Figure 11: Unlearning layer analysis $C_{f}$=15 for all layers [foster2024fast, golatkar2020eternal], last 1 layer and last 7 layers.
+TABLE IX: Large-scale Federated Unlearning Performance of Language and Graph Models on Yahoo! Answers
+Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
+Re-training FIUn
+$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
+Bert-Base-Cased
+1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.99 99.99 99.99 105.63 215.14 213.75
+$AD_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.99 0.00 0.00 1.56 2.84 2.79
+2 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.99 99.99 99.99 93.15 186.43 190.32
+$AD_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 15.13 0.00 10.21 1.51 2.75 2.81
+4 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.99 99.99 99.99 70.31 143.53 145.17
+$AD_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 26.73 0.00 6.41 1.90 2.91 3.14
+GCNN
+1 $AD_{r}\uparrow$ 99.56 99.93 96.43 99.69 99.99 97.64 91.54 68.69 95.67 33.95 67.31 101.64
+$AD_{f}\downarrow$ 98.81 99.99 94.64 0.00 0.00 0.00 0.00 0.00 9.52 0.42 0.77 0.76
+2 $AD_{r}\uparrow$ 99.56 99.93 96.43 99.99 99.99 98.31 63.89 64.29 71.64 28.60 57.14 86.41
+$AD_{f}\downarrow$ 99.11 99.99 95.60 0.00 0.00 0.00 0.00 0.00 0.00 0.44 0.87 0.86
+4 $AD_{r}\uparrow$ 99.56 99.93 96.43 99.99 99.99 98.34 93.17 63.07 97.41 21.63 42.60 64.31
+$AD_{f}\downarrow$ 99.07 99.99 95.67 0.00 0.00 0.00 0.00 0.00 25.37 0.34 0.79 0.84
+(a) Throughput, FL, TinyImageNet, ResNet18
+(b) Latency, FL, TinyImageNet, ResNet18
+(c) Throughput, FL, TinyImageNet, DenseNet161
+(d) Latency, FL, TinyImageNet, DenseNet161
+(e) Throughput, FL, Yahoo, Bert
+(f) Latency, FL, Yahoo, Bert
+(g) Throughput, FL, Yahoo, GCNN
+(h) Latency, FL, Yahoo, GCNN
+Figure 12: Scalability results of the FIUn method in FL.
+Settings: unlearning requests from 1,000 to 10,000 (step 1,000), 8,000 concurrent requests, DAG depth 3, models include TinyImageNet (ResNet18, DenseNet161) and Yahoo (Bert, GCNN).
+Line charts show throughput vs.
+transactions; bar charts show latency.
+VII-B Merged Label Unlearning Analysis
+This experiment assesses the applicability of the MFIM function to models with various distributions of unlearned labels.
+We analyze the unlearning effectiveness of various models across CIFAR100 DDPL.
+As demonstrated in Table VII, the average $AD_{r}$ is 89.72%, and the average $AD_{f}$ is 5.62%, confirming the unlearning effectiveness of our method.
+We evaluate unlearning on TinyImageNet under FL.
+Table VIII shows that FIUn achieves complete forgetting of target classes while keeping retained-class accuracy high (92.4%), demonstrating strong stability and generalization via the MFIM function.
+FIUn also greatly reduces cumulative unlearning time compared to retraining and maintains consistently low unlearning time across label distributions.
+Additional experiments are provided in Appendix B.
+We evaluate the MFIM function’s ability to manage merged label unlearning tasks within FL.
+As shown in Fig.
+8, our method achieves an average 88.54% performance improvement in AlexNet and DenseNet161 models over direct FIM use.
+In the ResNet18 model, its unlearning effectiveness is comparable to that of FIM.
+This is because MFIM reduces cumulative errors (caused by the manual threshold rule) by merging FIMs from all relevant roots into a single calculation, allowing parameter adjustments to be made in one step instead of applying the threshold rule to each root [foster2024fast, golatkar2020eternal], leading to more accurate and stable unlearning outcomes.
+In experiments with BERT-base-cased and GCNN on the Yahoo!
+Answers dataset within FL (Table IX), we tested 50 clients for multi-label unlearning.
+The $AD_{f}$ metric showed small increases in some cases (15.13% and 26.73%), but overall remained low.
+For GCNN, under $C_{f}$=2 and 4, $AD_{r}$ exceeded 63%, though FIUn slightly reduced retained accuracy due to the large, unoptimized FIM hyperparameter space.
+Nevertheless, FIUn was highly time-efficient, significantly faster than re-training, and achieved effective unlearning with only minor accuracy trade-offs.
+Takeaway–capable of multi-root unlearning.
+With up to 88.54% improvement, the MFIM function enables FIUn to consistently manage merged unlearning across different tasks and models, a capability that standard FIM approaches cannot achieve due to their inability to aggregate multiple FIMs into a single and unified FIM that encapsulates all root updates.
+VII-C Inherited Depth of Model Analysis
+We assess FIUn’s performance impact as inheritance depth increases in FL and IL, using consistent hyperparameters across all experiments.
+  • $\bullet$
     Federated Learning.
-In the FL framework, the CIFAR100 and TinyImageNet datasets are randomly divided into five segments.
-Each model trains using one segment of the data, adopting a binary tree structure as depicted in Figure 6.
+(Cyan in Fig.
+5) In FL, the CIFAR100 and TinyImageNet datasets are randomly divided into five segments.
+Each model trains using one segment of the data that adopts a binary tree structure.
 We evaluate whether our method can effectively unlearn specified labels as the depth of the binary tree increases.
+  • $\bullet$
     Incremental Learning.
-In the IL framework, as demonstrated in Figure 6, the starting model $w_{g}$ contains data from 170 classes.
-With each incremental learning step, one class of label data is added to assess whether our method can effectively unlearn specified labels.
-As shown in Figure 8 (i.e., single-class unlearning, see Appendix .4), we observe that under both frameworks, various datasets, and models with different unlearned labels, the performance of the proposed algorithm remains consistent regardless of the increased inherited depth.
-This consistency across both frameworks and models supports the notion that the impact of data on model parameters remains stable at fixed positions.
-Such stability ensures that our method maintains high performance even as the depth increases.
-6.4. Analysis of Parameter Layer Selection in Models
-We analyze how the number of model layers used to calculate the FIM impacts the performance of our algorithm.
-We conduct experiments within both FL and IL frameworks, with 15 label categories selected as unlearned labels.
-The DAG diagrams for these frameworks are illustrated in Figures 6 and 6.
-We calculate the FIM using the last layer, the last four layers, the last seven layers, and all layers to assess unlearning performance.
-As depicted in Figure 9, we observe that an increase in the number of layers used leads to diminishing unlearning performance across various datasets and models.
-This decline is attributable to the increase in the layer parameters, which inversely affects the $AD_{r}$ and $AD_{f}$ metrics.
-Moreover, a larger number of layers prolongs the time needed to compute the FIM.
-This finding underscores the effectiveness of our method, which relies on calculations from only the last layer.
-7. Related Work
-Although FL (Li et al., 2020b), DDPL (Li et al., 2020c), IL (Van de Ven et al., 2022), and TL (Weiss et al., 2016) have each made significant strides in addressing specific challenges, such as data privacy protection (Li et al., 2021), model training efficiency (Dean et al., 2012), knowledge retention and unlearning (Han et al., 2023), and cross-domain knowledge transfer (Rajapaksha et al., 2023), systematic exploration that combines these learning frameworks remains sparse.
-Most studies have merely attempted simple combinations of a few methods (Dai and Meng, 2023; Yoon et al., 2021; Dai et al., 2019; Miao et al., 2023) without touching the core of their commonality–model inheritance.
-This characteristic plays a crucial role in environments where unlearning is conducted on a large-scale.
-As data and models grow, the computational overhead increases exponentially with the depth of inheritance, while time consumption grows linearly with the depth, even when unlearning can be executed simultaneously at the same depth starting from the unlearning origin.
-This holds true regardless of the existing unlearning methods utilized, which are respectively discussed in the following.
-Federated Unlearning.
-FedEraser (Liu et al., 2020) used historical parameter updates from the central server to reconstruct the deleted model.
-Wu et al.
-(Wu et al., 2022a) subtracted historical updates and used knowledge distillation to restore model performance without client data.
-FRU (Yuan et al., 2023) recalled and calibrated historical updates to eliminate user contributions and accelerate federated recommendation reconstruction.
-These methods primarily relied on recalling and calibrating historical parameter updates of the model to achieve deletion or re-training.
-Halimi et al.
-(Halimi et al., 2022) reversed the learning process by maximizing local empirical loss and used PGD to perform deletion on the client side, while Wu et al.
-(Wu et al., 2022b) used class, client, and sample decoupling learning combined with Stochastic Gradient Ascent (SGA) and Elastic Weight Consolidation (EWC) to achieve deletion.
-These methods achieved deletion by optimizing the model’s loss function or gradients.
-Liu et al.
-(Liu et al., 2022) utilized first-order Taylor expansion and a diagonal empirical Fisher information matrix for accelerated re-training.
-Wang et al.
-(Wang et al., 2022) applied CNN channel pruning to remove specific category information, while FFMU (Che et al., 2023) employed nonlinear function analysis to enhance model deletion processes.
-These approaches streamline deletion or re-training by altering the model structure or employing targeted techniques.
-Additionally, KNOT (Su and Li, 2023) optimized re-training by introducing clustering aggregation and framing the client clustering issue as a dictionary minimization problem.
-Distributed Data-Parallel Unlearning.
-In DDPL knowledge unlearning, handling the unlearning process is restricted to the data shard on the specific device, followed by model aggregation and iterative training.
-If unlearning is required from a specific shard, only the model trained on that shard is re-trained.
-Approximate machine unlearning directly modifies trained model parameters, employing methods such as 1) Gradient-based approaches (Fan et al., 2023; Golatkar et al., 2020; Neel et al., 2021), which update model gradients to gradually diminish the data’s influence, and 2) Data perturbation (Tarun et al., 2023; Goel et al., 2024), which introduces noise to reduce specific data points’ impact on the model.
-These methods facilitate effective knowledge unlearning in DDPL.
-Incremental Unlearning.
-Research on unlearning on IL is still nascent, yet the Projected-Gradient Unlearning method (PGU) (Hoang et al., 2024) shows promise.
-PGU computes the gradients for the unlearning dataset and projects these onto the Core Gradient Space (CGS) derived from the training set.
-It then subtracts this projected component from the original gradients, positioning the residuals in a space orthogonal to the CGS, and updates the model parameters with these adjusted gradients.
-Effective for one-time data deletions and batch-wise unlearning, PGU is well-suited for ongoing incremental learning processes.
-Transfer Unlearning.
-Research on unlearning specific target task knowledge has been relatively scarce in TL.
-To address this challenge, (Sepahvand et al., 2024) proposed fine-tuning pre-trained models through data selection.
-Subsequently, the entire network was fine-tuned on the target task using gradient descent.
-What Sets Our Solution Apart.
-The above-mentioned existing studies on unlearning methods typically endure high computational complexity, significant resource demands, and implementation difficulties.
-These methods are time-consuming and resource-intensive, especially when deployed in large-scale environments, as they require accurate tracking and sequential updating of numerous model parameters.
-By contrast, the proposed FIUn method fully parallelizes unlearning, significantly enhancing computational efficiency, resource utilization, and adaptability.
-It effectively addresses the challenges faced by existing unlearning methods in inheritance-oriented model networks.
-8. Conclusion
-In this paper, we developed a new parallel unlearning framework for models with dependence and inheritance.
-To do this, we introduced the UMIG, which can effectively abstract model inheritance, e.g., in FL, DDPL, IL, and TL, using a DAG topology.
-We also developed the FIUn method and the merging FIM function, which can quickly locate impacted models and conduct parallel unlearning through the UMIG.
-Experimental results on various models and datasets showed that our method achieves complete unlearning for single-class labels, while maintaining an average accuracy of 94.53% for retained labels.
-For multi-class labels, the unlearning accuracy is only 1.07%, with retained label accuracy at 84.77%.
-Our method is 99% faster than existing unlearning methods, efficiently handling model updates regardless of inheritance depth.
+(Blue in Fig.
+5) In IL, the starting model $w_{g}$ includes data from 170 classes.
+With each learning step, a new class is added to evaluate the effectiveness of our method in unlearning specific labels.
+Figs.
+9 and 10 evaluate inheritance depth under FL and IL for $C_{f}\in{1,10}$.
+Overall, our method remains stable across depths; in FL, the unlearning accuracy $\bigtriangleup_{acc}$ stays near 100% even at large depths.
+In IL, especially TinyImageNet + DenseNet161, $\bigtriangleup_{acc}$ declines and fluctuates as depth grows, likely because DenseNet161’s dense connectivity accumulates class-specific features that are harder to erase, and TinyImageNet’s strong inter-class overlap amplifies category and parameter drift; depth-dependent adaptation and data stochasticity further contribute to variability.
+Takeaway–perfectly matching inheritance networks.
+FIUn exhibits the Hyper-Distance property, effectively unlearning across up to 50 levels of inheritance depth in various models and datasets, enabling simultaneous updates of all inherited models during unlearning tasks.
+VII-D Parameter Layer Selection in Models
+We evaluate the impact of the number of model layers used for FIM calculation on algorithm performance.
+Experiments were conducted in FL and IL frameworks with 15 label categories selected for unlearning.
+Fig.
+5 illustrates the DAG for these frameworks with Yellow and Blue models ($w_{g}$, $w_{a}$, $w_{b}$).
+FIM calculations were conducted on the last layer, last seven layers, and all layers using 10 random parameter sets.
+We apply the methods from [foster2024fast, golatkar2020eternal] (using all layers) for unlearning in inherited models.
+As shown in Fig.
+11, computing FIM on all layers leads to high variance in $\bigtriangleup_{acc}$.
+In contrast, the “Last 1” and “Last 7” settings show much smaller variance and achieve over 100% better unlearning.
+This suggests that focusing on partial layers—especially the last layer—yields more stable and consistent results.
+The reason is that early layers capture general features, while later layers, particularly the last, encode task-specific features.
+Thus, targeting the last layer removes task-specific information more efficiently, which fits well with our MFIM design.
+Takeaway–focusing on the last layer is the best for efficiency and unlearning effectiveness.
+FIUn achieves more efficient and effective unlearning by strategically performing FIM calculations on the weights of the last layer of models that the calculation of the MFIM function can benefit from.
+VII-E Unlearning Speed Analysis
+We conduct experiments on the inherited model $w_{b}$ using AlexNet and ResNet models, as shown in Fig.
+5, while the unlearning speed is presented in Fig.
+13.
+The results show that the unlearning time progressively increases with the number of unlearning layers, the parameters for AlexNet are 409600, 18833,764, and 20495268 for 1, 7, and all layers, respectively, while for ResNet18, they are 51200, 4903524, and 11227812, respectively.
+The results confirm that our proposed FIUn method with the last layer effectively enhances the efficiency of the unlearning process.
+Figure 13: Unlearning speed of different layers of models. FIUn speed is based on the FL framework with CIFAR100.
+VII-F Scalability Experiments
+To validate the scalability of FIUn, we conduct parallel experiments with the following key configurations:
+  • Unlearned data volume: Requests range from 1,000 to 10,000, in increments of 1,000.
+  • Concurrent requests: A concurrency level of 8,000 is maintained.
+  • Distributed participants: Experiments use a DAG depth of 3, with scalability tested on federated setups across models such as TinyImageNet (ResNet18, DenseNet161) and Yahoo (Bert, GCNN).
+Fig.
+12 reports FL scalability with FIUn, measuring throughput and latency on TinyImageNet (ResNet18, DenseNet161) and Yahoo (BERT, GCNN).
+FIUn shows near-linear throughput growth and stable latency as transaction volume increases, reflecting its efficient parallel unlearning; the setup is reproducible and indicates real-world applicability.
+Additional experiments appear in Appendix C.
+VIII Conclusion
+We developed a novel parallel unlearning framework for models with dependence and inheritance, utilizing a chronological DAG to abstract and visualize model inheritance in FL, DDPL, IL, and TL.
+The FIUn method and MFIM function efficiently identified affected models and enabled parallel unlearning through the DAG.
+Experiments across various models and datasets demonstrated that FIUn achieved complete unlearning for single-class labels while maintaining an average retained accuracy of 94.53%.
+For multi-class labels, unlearning accuracy was just 1.07%, with retained accuracy at 84.77%.
+Our method is 99% faster than existing ones and handles model updates efficiently at all inheritance depths.
+TABLE X: Federated Unlearning Performance on CIFAR100
+Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
+Re-training FIUn
+$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
+AlexNet
+1 $AD_{r}\uparrow$ 99.99 99.99 99.81 99.99 99.99 98.80 97.13 96.41 97.12 29.70 59.06 89.62
+${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00 0.09 0.11 0.13
+2 $AD_{r}\uparrow$ 99.99 99.99 99.86 99.99 99.99 99.51 91.89 90.78 95.66 28.69 57.81 84.97
+${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00 0.11 0.16 0.13
+4 $AD_{r}\uparrow$ 99.99 99.99 99.96 99.99 99.99 99.16 85.79 78.54 89.40 26.39 53.10 81.96
+${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00 0.10 0.16 0.13
+ResNet18
+1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 97.09 97.13 96.93 30.36 61.40 90.94
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.68 1.00 0.99
+2 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 96.66 98.10 97.44 28.30 57.39 84.06
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.69 1.03 1.09
+4 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 97.17 99.45 99.19 26.39 52.69 81.39
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.75 1.09 1.06
+TABLE XI: Incremental Unlearning Performance on CIFAR100
+Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
+Re-training FIUn
+$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
+AlexNet
+1 $AD_{r}\uparrow$ 99.99 99.99 99.81 99.99 99.99 98.80 98.97 94.39 94.98 22.95 43.83 63.45
+${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00 0.09 0.39 0.39
+2 $AD_{r}\uparrow$ 99.99 99.99 99.80 99.99 99.99 99.51 92.93 71.58 69.14 21.35 41.58 62.10
+${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00 0.11 0.27 0.26
+4 $AD_{r}\uparrow$ 99.99 99.99 99.96 99.99 99.99 99.16 88.77 59.62 58.81 20.52 41.12 60.69
+${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00 0.12 0.28 0.29
+ResNet18
+1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 91.71 90.05 90.82 26.35 47.12 66.93
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.30 0.80 0.85
+2 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 91.49 89.02 87.64 25.37 45.93 65.89
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.30 0.84 0.85
+4 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 84.71 83.42 84.54 24.83 43.23 65.38
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.31 0.86 0.87
+TABLE XII: Transfer Unlearning Performance on TinyImageNet
+Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
+Re-training FIUn
+$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
+DenseNet161
+1 $AD_{r}\uparrow$ 86.33 91.24 89.34 97.75 97.52 96.75 71.49 69.82 66.73 987.94 1986.35 1985.16
+${AD}_{f}\downarrow$ 94.50 94.50 96.70 0.00 0.00 0.00 0.00 0.00 0.00 2.41 4.15 4.34
+4 $AD_{r}\uparrow$ 86.33 91.24 89.34 96.42 96.41 95.13 64.24 55.32 51.18 1001.34 1994.53 1986.61
+${AD}_{f}\downarrow$ 89.12 91.23 91.84 0.00 0.00 0.00 0.00 0.00 0.00 2.34 4.71 4.26
+6 $AD_{r}\uparrow$ 86.33 91.24 89.34 95.72 95.43 93.58 58.69 47.38 42.57 1018.14 2008.89 2009.31
+${AD}_{f}\downarrow$ 86.57 89.25 90.28 0.00 0.00 0.00 0.00 0.00 0.00 2.16 4.64 4.46
+ResNet18
+1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 98.66 98.07 96.74 142.15 243.75 241.61
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.42 0.75 0.78
+4 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 98.84 98.38 97.38 143.46 244.34 245.74
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.42 0.74 0.75
+6 $AD_{r}\uparrow$ 99.96 99.99 99.99 99.99 99.99 99.99 98.82 98.91 98.98 147.27 245.35 246.41
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.56 0.89 0.74
+TABLE XIII: Distributed Data-Parallel Unlearning Performance on TinyImageNet
+Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
+Re-training FIUn
+$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
+DenseNet161
+1 $AD_{r}\uparrow$ 99.99 87.16 85.21 99.99 89.39 87.24 91.82 86.63 83.58 980.53 988.42 991.75
+${AD}_{f}\downarrow$ 99.99 92.75 90.41 0.00 0.00 0.00 0.00 0.00 0.00 4.34 2.61 2.13
+4 $AD_{r}\uparrow$ 99.96 87.00 85.21 99.96 88.35 89.34 87.42 86.56 84.37 1004.63 1006.32 1005.332
+${AD}_{f}\downarrow$ 99.96 86.97 88.85 0.00 0.00 0.00 0.00 0.00 0.00 4.31 2.05 2.35
+6 $AD_{r}\uparrow$ 99.99 87.00 85.21 99.96 87.96 88.15 82.50 86.82 84.47 1024.25 1036.14 1026.74
+${AD}_{f}\downarrow$ 99.99 89.35 86.08 0.00 0.00 0.00 0.00 0.00 0.00 4.16 2.24 2.21
+ResNet18
+1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.19 99.99 99.99 112.52 117.34 114.52
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 3.56 1.67 1.47
+4 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.99 99.99 99.99 109.32 108.74 110.53
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.32 3.54 1.74 1.37
+6 $AD_{r}\uparrow$ 99.96 99.99 99.99 99.96 99.96 99.96 99.38 99.99 99.99 107.78 108.32 107.17
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 3.53 1.84 1.64
+Appendix A General Label Unlearning Analysis
+Table X to Table XIII show that, on CIFAR100, FIUn achieves strong performance under federated unlearning and incremental unlearning frameworks.
+Similarly, on TinyImageNet, FIUn exhibits excellent performance under transfer unlearning and distributed data-parallel unlearning.
+Re-training on TinyImageNet takes a significantly longer time than CIFAR100.
+In comparison, FIUn maintains a high unlearning accuracy while keeping the unlearning time extremely short.
+Appendix B Merged Label Unlearning Analysis
+Tables XIV to XVII show that, in the federated unlearning setting on CIFAR100, our method achieves good performance on multi-label distributions.
+On TinyImageNet, under the distributed data-parallel unlearning setting, it demonstrates excellent performance on multi-label distributions.
+For multiple label distributions, our proposed MFIM function accelerates the unlearning speed of inherited models while maintaining high unlearning accuracy.
+Appendix C More Datasets and Large Scale Analysis
+Tables XV and XVI present the experimental results of unlearning using a ViT model on ImageNet and TinyImageNet datasets.
+Table XV shows that for ImageNet, when unlearning one class ($C_{f}=1$), the accuracy on retained data ($AD_{r}$) remains stable (73.8% $\rightarrow$ 73.64%) while the accuracy on unlearned data ($AD_{f}$) drops from 100% to 0%, with an unlearning time of 2.1s.
+For ten classes ($C_{f}=10$), $AD_{f}$ decreases from 89.04% to 0%, with $AD_{r}$ slightly reduces to 73.1% and an unlearning time of 4.33s.
+In the TinyImageNet experiment, as shown in Table XVI, we simulate a federated learning setup with 100 clients, each running the same ViT model.
+For both $C_{f}=1$ and $C_{f}=10$, $AD_{f}$ reaches 0% after unlearning, while $AD_{r}$ remains above 95%, with unlearning times ranging from 3.79s to 13.14s across clients.
+TABLE XIV: Federated Unlearning Performance on CIFAR-100 with Multiple Label Distribution
+Model $\cap_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
+Re-training FIUn
+$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
+AlexNet
+60% $AD_{r}\uparrow$ 99.99 99.99 99.81 98.80 99.99 99.99 69.42 81.68 75.05 59.26 30.13 29.70
+${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00 0.15 0.08 0.09
+40% $AD_{r}\uparrow$ 99.99 99.99 99.80 99.51 99.99 99.99 61.66 79.77 72.96 53.84 29.58 28.69
+${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00 0.11 0.09 0.19
+20% $AD_{r}\uparrow$ 99.99 99.99 99.96 99.16 99.99 99.99 59.57 77.47 77.97 57.36 27.48 26.39
+${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.12 0.00 0.00 0.10 0.07 0.18
+ResNet18
+60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 93.15 98.76 97.81 67.31 33.18 30.36
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.60 0.32 0.29
+40% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 93.78 99.04 98.98 52.33 28.13 28.30
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.59 0.34 0.30
+20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 88.54 98.79 98.62 56.61 23.37 26.39
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.58 0.28 0.30
+TABLE XV: ImageNet Experiment Results for Unlearning with Pretrained ViT Model
+Unlearning Label AD_r (%) AD_f (%) AD_r after Unlearning (%) AD_f after Unlearning (%) Unlearning Time (s)
+One Class ($C_{f}=1$) 73.8 100 73.64 0 2.1
+Ten Classes ($C_{f}=10$) 73.8 89.04 73.1 0 4.33
+TABLE XVI: TinyImageNet Experiment Results for Unlearning in Federated Learning Setup with ViT Model
+Unlearning Case Client Unlearning Time (s) $AD_{f}$ after Unlearning (%) $AD_{r}$ after Unlearning (%)
+One Class ($C_{f}=1$) Client 1 3.79 0.00 95.63
+Client 2 7.94 0.00 99.98
+Client 3 12.45 0.00 99.63
+Ten Classes ($C_{f}=10$) Client 1 3.84 0.00 99.99
+Client 2 7.96 0.00 99.99
+Client 3 13.14 0.00 99.99
+TABLE XVII: Distributed Data-Parallel Unlearning Performance on TinyImageNet with Multiple Label Distribution
+Model $\cap_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Unlearning Time (s)
+Re-training FIUn
+$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
+DenseNet161
+60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 76.86 89.47 82.53 994.32 998.46 1000.66
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 4.69 2.39 2.30
+40% $AD_{r}\uparrow$ 99.99 99.99 99.50 99.99 99.99 99.99 74.03 89.47 83.18 1010.45 1011.14 1009.53
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 4.12 2.69 2.14
+20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 73.48 89.13 83.78 1021.34 1020.67 1022.64
+${AD}_{f}\downarrow$ 99.99 99.99 99.91 0.00 0.00 0.00 0.00 0.00 0.00 4.98 2.64 2.59
+ResNet18
+60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 95.16 95.92 98.86 123.74 126.43 149.32
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 3.16 1.65 1.68
+40% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 94.43 95.92 98.27 128.43 127.64 155.32
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 3.97 1.46 1.59
+20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 94.71 95.92 98.48 129.35 126.43 154.83
+${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 3.89 1.54 1.38